### v2

# M-estimation for Gaussian processes with time-inhomogeneous drifts from high-frequency data

Abstract
We propose a contrast-based estimation method for Gaussian processes with time-inhomogeneous drifts, observed under high-frequency sampling. The process is modeled as the sum of a deterministic drift function and a stationary Gaussian component with a parametric kernel. Our method constructs a local contrast function from adjacent increments, which avoids inversion of large covariance matrices and allows for efficient computation. We prove consistency and asymptotic normality of the resulting estimators under general ergodicity conditions. A distinctive feature of our approach is that the drift estimator attains a nonstandard convergence rate, stemming from the direct Riemann integrability of the drift density. This highlights a fundamental difference from standard estimation regimes. Furthermore, when the local contrast fails to identify all parameters in the covariance kernel, moment-based corrections can be incorporated to recover identifiability. The proposed framework is simple, flexible, and particularly well suited for high-frequency inference with time-inhomogeneous structure.
Keywords: Gaussian processes, high-frequency data, time-inhomogeneous drift, contrast-based estimation, method of moments
MSC2010: 62M10; 62F12, 60G15.
1 Introduction
Gaussian processes (GPs) are powerful tools for probabilistic modeling and forecasting of time-evolving phenomena. Their nonparametric nature and built-in uncertainty quantification make them well-suited for modeling irregular or noisy data without assuming a specific functional form. For these reasons, GPs are widely used in applications ranging from machine learning to time series analysis.
A central task in Gaussian process modeling is the estimation of structural parameters from discrete observations. In particular, time series data often exhibit both deterministic long-term trends and stochastic short-term fluctuations. This motivates models that combine a time-inhomogeneous drift with a stationary Gaussian process. Such models naturally arise in settings such as mortality forecasting and environmental statistics, where gradual trends are superimposed with random variation.
We consider the following stochastic process $X=(X_{t})_{t\geq 0}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$:
$X_{t}=Z_{t}+\int_{0}^{t}\mu(s)\,\mathrm{d}s,\quad t>0,$
where $\mu:[0,\infty)\to\mathbb{R}$ is a deterministic function, and $Z=(Z_{t})_{t\geq 0}$ is a Gaussian process defined as follows.
Definition 1.1.
A stochastic process $Z=(Z_{t})_{t\geq 0}$ is called a Gaussian process if and only if any finite-dimensional distribution is multivariate Gaussian: for any $t_{1},\dots,t_{d}>0$, there exist a mean vector $m\in\mathbb{R}^{d}$ and a positive definite $d\times d$ matrix $\Sigma$ such that
$(Z_{t_{1}},Z_{t_{2}},\dots,Z_{t_{d}})\sim N_{d}(m,\Sigma),$
where $N_{d}(m,\Sigma)$ denotes the $d$-dimensional normal distribution with mean $m$ and covariance matrix $\Sigma$.
Definition 1.2.
A Gaussian process $Z$ is said to be stationary if there exists a function $K:[0,\infty)\to\mathbb{R}$ such that
$\mathbb{E}[Z_{t}Z_{s}]=K(|t-s|),\quad t,s\geq 0.$
That is, the covariance between $Z_{t}$ and $Z_{s}$ depends only on their time lag. The function $K$ is called the kernel function. Moreover, $Z$ is said to be centered if the mean function $m(t):=\mathbb{E}[Z_{t}]$ is identically zero.
A stationary Gaussian process is fully characterized by its kernel function, which governs properties such as smoothness and temporal dependence. Accurate estimation of the kernel parameters is crucial for model performance, and has been widely studied; see Rasmussen and Williams [17] for a comprehensive reference.
Two of the most widely used approaches for parameter estimation in GP models are Maximum Likelihood Estimation (MLE) and Maximum A Posteriori (MAP) estimation. The MAP method incorporates prior information by maximizing the posterior density, while MLE relies solely on the marginal likelihood. Since $n$ observations from a GP follow an $n\times n$ multivariate normal distribution, both methods require construction and inversion of the $n\times n$ covariance matrix. These steps incur computational costs of $O(n^{2})$ and $O(n^{3})$, respectively, and become infeasible for large $n$ or high-frequency data. Furthermore, as noted by Karvonen et al. [10], MLE may suffer from ill-posedness in near-deterministic or noise-free settings.
To address the computational and statistical challenges of MLE in such regimes, a variety of approximate methods have been proposed. Minden et al. [15] developed recursive skeletonization factorizations to accelerate MLE for spatial GPs. Composite likelihood estimation (CLE), developed by Cox and Reid [4], Davis and Yau [5], and Varin et al. [21], constructs contrast functions based on low-dimensional marginals. Bennedsen et al. [2] apply CLE to high-frequency Gaussian processes by combining $q$-dimensional marginal densities (typically $q=3$), which avoids full matrix inversion but still requires repeated inversion of $q\times q$ covariance matrices.
An alternative class of methods uses frequency-domain approximations. Whittle [22] proposed likelihood in the frequency domain, now known as the Whittle likelihood, for stationary Gaussian processes, which reduces computational complexity to $O(n\log n)$. Building on this idea, Fukasawa and Takabatake [8] developed a high-frequency extension of the Whittle likelihood that achieves asymptotic efficiency for self-similar stationary Gaussian noise. Takabatake [19] further extended this framework to models driven by fractional Brownian motion with stochastic drift, constructing a quasi-Whittle likelihood estimator and proving consistency and asymptotic normality.
A different line of work, exemplified by Kobayashi et al. [12], focuses on rigorous continuous-time likelihood inference. They investigated the “exact MLE” of the drift parameter based on continuous-time likelihood, establishing the LAN property of the likelihood, and further analyzed its discretization under high-frequency sampling, proving asymptotic efficiency.
In contrast to the above methods, our approach originates from pseudo-likelihood techniques for stochastic differential equations (SDEs); see Kessler [11]. We construct a contrast-based estimator from local increments of the observed process. The contrast function depends only on scalar conditional variances and does not require matrix inversion. This yields simple estimating equations that are both computationally efficient and theoretically tractable under high-frequency sampling.
Interestingly, this setting also leads to a nonstandard convergence rate for the drift estimator, which arises from the direct Riemann integrability (DRI) inherent in the ergodic structure of the process. Moreover, our framework directly handles time-dependent drift functions, unlike most spectral-domain methods which require prior detrending.
The key distinction from CLE lies in the information structure. CLE aggregates multiple lagged values within each block, exploiting high-order correlations. Our method, by contrast, uses only second-order local differences. Although this minimal structure may lead to identifiability issues for some parameters in the kernel, such issues can be addressed by incorporating moment-based corrections.
We assume a smooth parametric kernel function to facilitate theoretical analysis. However, our method is also applicable to processes with nonsmooth kernels, such as the Ornstein-Uhlenbeck kernel, which is not differentiable at the origin. In such cases, the kernel can be approximated by a family of smooth kernels constructed via mollification. This mollified approximation enables the application of our framework to a broader class of Gaussian processes. See Subsection 4.2 for an illustrative example.
Beyond the specific estimators considered, the asymptotic theory developed in this paper—some limit theorems including uniform laws of large numbers under high-frequency sampling—offers fundamental tools for future methodological developments. In particular, these results may serve as a foundation for hybrid procedures combining time- and frequency-domain techniques, or for investigating the asymptotic efficiency of more general contrast-based methods.
Our method has the following advantages:
  • No matrix inversion is required: The contrast depends only on scalar variances, avoiding the computational burden of full MLE or CLE.
  • Theoretically justified under general ergodic Gaussian processes: Consistency and asymptotic normality are established without assuming Markovianity or spectral representation.
  • High-frequency suitability: Designed for dense-sampling regimes, the method is numerically stable and scalable.
  • Moment-based correction for identifiability: When local contrast fails to identify kernel parameters, moment estimators can restore identifiability.
  • Applicability to nonsmooth kernels: Even when the kernel function is not differentiable at the origin (e.g., the Ornstein–Uhlenbeck kernel), our method remains applicable by approximating it via a family of smooth kernels using mollifiers.
  • Potential extensibility to frequency-domain methods: The estimated drift can be subtracted to allow for Whittle-based inference on residuals, enabling hybrid approaches for long-term structure. Importantly, our asymptotic results—such as the uniform law of large numbers, central limit theorems type results, under high-frequency sampling—provide essential theoretical tools for developing such two-step procedures in the future.
  • Solid theoretical foundation for future extensions: The asymptotic results established in this work—such as uniform laws of large numbers and central limit theorems under high-frequency sampling—provide essential tools for analyzing more complex procedures, including hybrid methods or efficiency theory for general contrast-based estimators.
The paper is organized as follows. Section 2 gives the Gaussian process model with time-inhomogeneous drift and describes the basic estimation framework. Section 3 develops the asymptotic theory for both contrast-based and moment-type $M$-estimators, including consistency and asymptotic normality. Section 4 provides concrete examples and analytical forms of the estimators.Section 6 contains the main proofs of the asymptotic results. Section 5 concludes the paper with additional remarks and potential extensions. Appendix A summarizes some limit theorems for ergodic Gaussian processes. Appendix B collects technical lemmas and auxiliary results used in the proofs, which are analogous to those in Appendix A but adapted to drifted data. Appendix C provides supplementary discussion on the ergodicity of Gaussian processes.
Notation
  • The random vector $Z$ follows the normal distribution with mean vector $m$ and covariance matrix $\Sigma$, we write $Z\sim{\cal N}(m,\Sigma)$.
  • The probability density function of ${\cal N}(0,\Sigma)$ is given by $\phi_{\Sigma}$.
  • For a centered stationary Gaussian process $Z=(Z_{t})_{t\geq 0}$ with the kernel function $K$, we write $Z\sim GP(0,K)$.
  • For a subet $\subset\mathbb{R}^{p}, S¯\overline{S}$ is the closure of $S$ w.r.t. the Euclidian norm.
  • For a function $f(x,y):\mathbb{R}^{d}\times\mathbb{R}^{d^{\prime}}\to\mathbb{R}$ and $x=(x_{1},\dots,x_{d})^{\top}$,
    $\partial_{x}f:=\frac{\partial f}{\partial x}=\left(\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{d}}\right)^{\top}\in\mathbb{R}^{d},\quad\partial_{x}^{2}f:=\partial_{x}\partial_{x}^{\top}f=\begin{pmatrix}\frac{\partial^{2}f}{\partial x_{1}\partial x_{1}}&\dots&\frac{\partial^{2}f}{\partial x_{1}\partial x_{d}}\\ \vdots&\ddots&\vdots\\ \frac{\partial^{2}f}{\partial x_{d}\partial x_{1}}&\dots&\frac{\partial^{2}f}{\partial x_{d}\partial x_{d}}\end{pmatrix}\in\mathbb{R}^{d\times d}.$
    if the partial derivatives exist, where $\top$ stands for the transpose.
2 Models and Assumptions
2.1 Gaussian processes with time-inhomogeneous drifts
Consider a stochastic process driven by a Gaussian process $Z$:
$X_{t}=Z_{t}+\int_{0}^{t}\mu(s)\,\mathrm{d}s,\quad X_{0}=Z_{0}$ (2.1)
where $\mu:[0,\infty)\to\mathbb{R}$ and $Z=(Z_{t})_{t\geq 0}\sim GP(0,K)$, a centered stationary Gaussian process with the kernel function $K$. The goal of the paper is to estimate the mean density $\mu$ and kernel functions $K$ from discrete samples of $X$ as follows:
$X_{t_{0}},\ X_{t_{1}},\ \dots,\ X_{t_{n-1}},\ X_{t_{n}},$
where $t_{i}=ih_{n}$ with $h_{n}>0$ for $i=1,2,\dots,n$. In asymptotic theory, we assume high-frequency sampling over a long time horizon, which is standard in modern applications where sufficiently dense observations are available:
$h_{n}\to 0,\quad nh_{n}\to\infty,\quad\mbox{as $n\to\infty$. }$ (2.2)
For that purpose, we consider parametric families for $\mu$ and $K$ as follows:
$\left\{\mu_{\xi}:[0,\infty)\to\mathbb{R}\,|\,\xi\in\overline{\Xi}\right\};\quad\left\{K_{\sigma}:[0,\infty)\to(0,\infty)\,|\,\sigma\in\overline{\Pi}\right\},$
where $\Xi\subset\mathbb{R}^{p}$ and $\Pi\subset\mathbb{R}^{q}$ are open bounded subset, and set $\Theta:=\Xi\times\Pi$. We suppose that there is the true values of parameters:
$\theta_{0}=(\xi_{0},\sigma_{0})\in\Theta;\quad\mu_{\xi_{0}}\equiv\mu;\quad K\equiv K_{\sigma_{0}}.$
A 1.
$K(t)\to 0$ as $t\to\infty$.
A 2.
$K\in C^{2}([0,\infty))$ with $\partial_{t}K(0)=0$ and $|\partial_{t}^{2}K(0)|>0$, that is,
$K(t)=K(0)+\frac{1}{2}\partial_{t}^{2}K(0)t^{2}+o(t^{3}),\quad t\to 0.$
A 3.
$K\in C^{2}([0,\infty))$ and
$\partial_{t}K(t)\to 0,\quad t\to\infty.$
Condition A1 is a mixing-type condition. In fact, the process $Z$ is weakly mixing under A1; see Maruyama [14]. This ensures the ergodicity of $Z$ in the sense of Corollary C.1.
This smoothness condition relates to the regularity of sample paths of the underlying Gaussian process. The assumption A2 is just the Taylor expansion and the remainder will be $o(t^{3})$ since $K$ is symmetric. However, such a smoothness is not always standard; in fact, certain important examples such as the Ornstein-Uhlenbeck (O-U) process (see Example 2.2) do not satisfy it. Nonetheless, consistent estimators of $\theta$ can be constructed by approximating the non-smooth kernel with a smooth mollified version $K_{\varepsilon}$ such that $K_{\varepsilon}\to K$ as $\varepsilon\to 0$; see Section 4.2.
We shall give some examples on stationary kernels.
Example 2.1.
Let us give some examples for kernel functions satisfying A1 and A2.
  • Gaussian kernel (Radial Basis Function):
    $K_{\sigma}(t)=\alpha\exp\left(-\frac{\beta}{2}t^{2}\right),\quad\sigma=(\alpha,\beta)\in\mathbb{R}^{2},$
    with
    $K_{\sigma}(t)=\alpha-\alpha\beta\,t^{2}+o(t^{3}),\quad t\to 0.$
  • Matérn kernel: (this is of $C^{2}$ as $\nu>2$)
    $K_{\sigma}(t)=\alpha\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}\beta|t|\right)^{\nu}B_{\nu}\left(\sqrt{2\nu}\beta|t|\right),\quad\sigma=(\alpha,\beta,\nu)\in\mathbb{R}^{3}$
    where $B_{\nu}$ is the modified Bessel function of 2nd kind. It is known that
    $K_{\sigma}(t)=\alpha-\alpha\beta^{2}\frac{\nu}{2\nu-1}t^{2}+o(t^{3}),\quad t\to 0.$
  • Rational Quadratic kernel
    $K_{\sigma}(t)=\alpha\left(1+\frac{\beta^{2}t^{2}}{2\gamma}\right)^{-\gamma},\quad\sigma=(\alpha,\beta,\gamma)\in\mathbb{R}_{+}^{3},$
    with
    $K_{\sigma}(t)=\alpha-\frac{\alpha\beta^{2}}{2}t^{2}+o(t^{3}),\quad t\to 0,$
    In particular, it is called ‘Cauchy kernel’ as $\alpha=1$.
Example 2.2.
Exponential (O-U) kernel The following kernel is also important in applications. It is called the exponentail kernel or the Ornstein-Uhlenbeck kernel:
$K_{\sigma}(t)=\alpha\exp\left(-\beta|t|\right),\quad\sigma=(\alpha,\beta)\in\mathbb{R}_{+}^{2},$
which is not smooth at $t=0$ since it includes $|t|$ in the exponent. However, we can approximate this kernel with a mollifier $\varphi\in C^{\infty}(\mathbb{R})$ such as: ${\rm supp}\,\varphi$ is compact;
$\int_{\mathbb{R}}\varphi(s)\,\mathrm{d}s=1;\quad\varphi_{\epsilon}(t):={\epsilon}^{-1}\varphi\left({\epsilon}^{-1}t\right)\to\delta_{0}(t)\quad{\epsilon}\to 0,$
where $\delta_{0}$ is Dirac’s delta function. Consider a smoothed kernel
$K_{\sigma}^{({\epsilon})}(t)=\int_{\mathbb{R}}K_{\sigma}(t-s)\varphi_{\epsilon}(s)\,\mathrm{d}s\to K_{\sigma}(t)\quad{\epsilon}\to 0.$
Then, it follows for $m\in\mathbb{N}$ that
$\partial_{t}^{m}K_{\sigma}^{({\epsilon})}(t):=\int_{\mathbb{R}}K_{\sigma}(t-s)\partial^{m}\varphi_{\epsilon}(s)\,\mathrm{d}s\to\partial_{t}^{m}K_{\sigma}(0),\quad{\epsilon}\to 0,$
where the last $\partial_{t}^{m}K_{\sigma}(0)$ is a ‘generalized’ derivative. Therefore, a smooth $K_{\sigma}^{({\epsilon})}$ is available as an approximation of non-smooth $K(t)$ in practice. For example, using the Laplace mollifier
$\varphi_{\epsilon}(s)=\frac{1}{2{\epsilon}}e^{-\frac{|s|}{{\epsilon}}},$
we have for each $t>0$:
$K_{\sigma}^{({\epsilon})}(t)$ $=\alpha e^{-\beta t}\int_{\mathbb{R}}e^{\beta s}\varphi_{\epsilon}(s)\mathbf{1}_{\{s\leq t\}}\,\mathrm{d}s+\alpha e^{\beta t}\int_{\mathbb{R}}e^{-\beta s}\varphi_{\epsilon}(s)\mathbf{1}_{\{s>t\}}\,\mathrm{d}s\to\frac{\alpha}{1+\beta{\epsilon}}\quad(t\to 0)$
$\partial_{t}K_{\sigma}^{({\epsilon})}(t)$ $=-\alpha\beta e^{-\beta t}\int_{\mathbb{R}}e^{\beta s}\varphi_{\epsilon}(s)\mathbf{1}_{\{s\leq t\}}\,\mathrm{d}s+\alpha\beta e^{\beta t}\int_{\mathbb{R}}e^{-\beta s}\varphi_{\epsilon}(s)\mathbf{1}_{\{s>t\}}\,\mathrm{d}s,$
$\to-\alpha\beta\int_{-\infty}^{0}e^{\beta s}\varphi_{\epsilon}(s)\,\mathrm{d}s+\alpha\beta\int_{0}^{\infty}e^{-\beta s}\varphi_{\epsilon}(s)\,\mathrm{d}s=0,\quad(t\to 0)$
$\partial_{t}^{2}K_{\sigma}^{({\epsilon})}(t)$ $=\alpha\beta^{2}\left[e^{-\beta t}\int_{\mathbb{R}}e^{\beta s}\varphi_{\epsilon}(s)\mathbf{1}_{\{s\leq t\}}\,\mathrm{d}s+e^{\beta t}\int_{\mathbb{R}}e^{-\beta s}\varphi_{\epsilon}(s)\mathbf{1}_{\{s>t\}}\,\mathrm{d}s\right]-2\alpha\beta\varphi_{\epsilon}(t)$
$\to\frac{\alpha\beta^{2}}{1+\beta{\epsilon}}-\frac{\alpha\beta}{{\epsilon}}.$
Moreover, as $t\downarrow 0$, using this explicit form we obtain the following expansion:
$K_{\sigma}^{({\epsilon})}(t)$ $=\frac{\alpha}{1+\beta{\epsilon}}-\left(\frac{\alpha\beta}{{\epsilon}}-\frac{\alpha\beta^{2}}{1+\beta{\epsilon}}\right)t^{2}+o(t^{3}),\quad(t\to 0).$
Hence, the mollified kernel $K_{\sigma}^{({\epsilon})}$ is smooth and satisfies the condition A2 for any fixed ${\epsilon}>0$.
That is, instead of modeling the data by an exact Ornstein–Uhlenbeck process, one can model it by a Gaussian process with kernel $K_{\sigma}^{(\varepsilon)}$ using a small $\varepsilon>0$. See the example in Subsection 4.2 for an illustration of this approach.
Remark 2.1.
As for the assumption A2, there exists a Gaussian process with $\partial_{t}^{2}K_{\sigma}(0)=0$. For example,
$K_{\sigma}(t)=\begin{cases}\exp\left(-\frac{1}{1-\left(t/\sigma\right)^{2}}\right)&\text{if }|t|<\sigma,\\ 0&\text{otherwise},\end{cases}$
is of $C^{\infty}$-class, and all the derivatives at $t=0$ is identically zero. Since we can also confirm the semi-positive definiteness of this kernel, this is a Gaussian kernel. However, such a Gaussian process with ‘flat’ derivatives at zero is impractical because the sample paths are too smooth due to the small quadratic variation. From an applied point of view, we are interested in processes with higher volatilities. Hence, A2 is not so strong assumption.
In what follows, we shall use the concept of directly Riemann integrable (DRI): A non-negative function $g:\mathbb{R}\to[0,\infty)$ is said to be DRI if its upper and lower Riemann sums over the whole real line converge to the same finite limit, as the mesh of the partition vanishes:
$\lim_{h\downarrow 0}\sum_{k\in\mathbb{Z}}\underline{g}_{k,h}\cdot h=\lim_{h\downarrow 0}\sum_{k\in\mathbb{Z}}\overline{g}_{k,h}\cdot h=:\int_{\mathbb{R}}g(s)\,\mathrm{d}s\in(-\infty,\infty),$
where $\overline{g}_{k,h}:=\sup_{z\in[kh,(k+1)h)}g(z)$ and $\underline{g}_{k,h}:=\inf_{z\in[kh,(k+1)h)}g(z)$. If the function $g$ may also take negative values, it is said to be DRI if both its positive and negative parts $g^{+}$ and $g^{-}$ are so. For more details, see, e.g., Asmussen [1], Section V.4; Feller [7], Section XI.1; Rolski et al. [18]; Caravenna [3]; and references therein.
Remark 2.2.
Note that if $g$ is DRI, then $g$ is bounded and continuous a.e. with respect to the Lebesgue measure; see Asmussen [1], Proposition V.4.1.
Remark 2.3.
In our sampling scheme (2.2), it follows for a DRI function $g:[0,\infty)\to\mathbb{R}$ that
$\lim_{n\to\infty}\sum_{i=1}^{n}g(t_{i-1})h_{n}=\int_{0}^{\infty}g(s)\,\mathrm{d}s,$ (2.3)
since, under $nh_{n}\to\infty$,
$\lim_{n\to\infty}\left[\sum_{i=1}^{\infty}g(t_{i-1})h_{n}-\sum_{i=1}^{n}g(t_{i-1})h_{n}\right]=\lim_{n\to\infty}\int_{nh_{n}}^{\infty}g(s)\,\mathrm{d}s\to 0.$
Therefore, when we use the convergence (2.3), the condition $nh_{n}\to\infty$ is always required.
We further impose the following assumptions on the parametric model, which will be introduced as needed in the discussion below.
B 1.
For any $\sigma\in\overline{\Pi}$, $K_{\sigma}\in C^{2}([0,\infty))$ with $\partial_{t}K_{\sigma}(0)=0$ and $|\partial_{t}^{2}K_{\sigma}(0)|\in(0,\infty)$, that is,
$\sup_{\sigma\in\overline{\Pi}}\left|\frac{K_{\sigma}(t)-K_{\sigma}(0)-\frac{1}{2}\partial_{t}^{2}K_{\sigma}(0)t^{2}}{t^{2}}\right|\to 0,\quad t\downarrow 0.$
B 2.
For any $\xi\in\Xi$, $\mu_{\xi}\in C^{1}(\mathbb{R})$ and $\sup_{t>0,\xi\in\overline{\Xi}}|\partial_{t}\mu_{\xi}(t)|<\infty$.
B 3.
For any $\xi\in\Xi$, $\mu_{\xi}$ is DRI on $\mathbb{R}$.
B 4.
$\mu_{\xi}=\mu_{\xi_{0}}\ a.e.$ for any $t\in[0,\infty)$ implies that $\xi=\xi_{0}$.
B 5.
$\partial^{2}_{t}K_{\sigma}(0)=\partial^{2}_{t}K_{\sigma_{0}}(0)$ implies that $\sigma=\sigma_{0}$.
B 6.
For any $\xi\in\Xi$, the function $t\mapsto\partial_{\xi}\mu_{\xi}(t)$ is DRI on $0,\infty)$.
B 7.
For any $\xi\in\Xi$, the function $t\mapsto\partial_{\xi}^{2}\mu_{\xi}(t)$ is bounded.
B 8.
There exists a bounded function $\gamma\in L^{1}([0,\infty))$ such that $\sup_{\xi\in\Xi}|\partial_{\xi}\mu_{\xi}(t)|\leq\gamma(t)$ for each $t\geq 0$.
3 M-estimation for sampled Gaussian processes
3.1 Local-Gauss contrast
We use the notation that, for a process $X=(X_{t})_{t\geq 0}$,
$\Delta_{i}^{n}X:=X_{t_{i}}-X_{t_{i-1}},$
the increment of $X$ on $(t_{i-1},t_{i}]\ (i=1,2,\dots,n)$. Note that
$\Delta_{i}^{n}X-\mu(t_{i-1}^{n})h_{n}\approx\Delta_{i}^{n}Z=(1,-1)\begin{pmatrix}Z_{t_{i}^{n}}\\ Z_{t_{i-1}^{n}}\end{pmatrix}\sim{\cal N}(0,2[K(0)-K(h_{n})]).$
Hence, we will use the following local-Gauss contrast function:
$\ell_{n}(\xi,\sigma)=\frac{1}{n}\sum_{i=1}^{n}\frac{\left(\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}\right)^{2}}{2[K_{\sigma}(0)-K_{\sigma}(h_{n})]}+\log\left(2h_{n}^{-2}[K_{\sigma}(0)-K_{\sigma}(h_{n})]\right).$ (3.1)
Note that $2h_{n}^{-2}$ in the logarithm of the second term is for an appropriate scaling to obtain a proper limt of $\ell_{n}(\theta)$.
We consider a minimum contrast estimator ($M$-estimator) defined as follows:
$\widehat{\theta}_{n}:=(\widehat{\xi}_{n},\widehat{\sigma}_{n})=\arg\min_{\theta\in\overline{\Theta}}\ell_{n}(\theta).$ (3.2)
Remark 3.1.
To estimate $\xi_{0}$, we can optimize the following simplified estimating function:
$\widehat{\xi}_{n}=\arg\min_{\xi\in\Xi}\sum_{i=1}^{n}\left(\Delta_{i}^{n}X-h_{n}\mu_{\xi}(t_{i-1})\right)^{2},$
which corresponds to the least squares estimation, and it does not require an estimator for $\sigma_{0}$. This often yields an explicit form for $\widehat{\xi}_{n}$; see, e.g., Example 4.1.
Remark 3.2 (Relation to composite likelihood methods).
Our contrast function is constructed from the sequence of increments $\Delta_{i}^{n}X:=X_{t_{i}}-X_{t_{i-1}}$, based on the assumption that these are approximately centered Gaussian with variance $K_{\sigma}(0)-K_{\sigma}(h_{n})$. This yields a pseudo-likelihood that is computationally efficient, as it avoids inversion of covariance matrices. This approach shares a structural similarity with the composite likelihood estimation (CLE) method proposed by Bennedsen et al. [2], which constructs a contrast by aggregating marginal Gaussian likelihoods over $q$-dimensional vectors. In particular, when $q=2$, CLE uses the full bivariate Gaussian likelihood of $(X_{t_{i-1}},X_{t_{i}})^{\top}$, incorporating the full $2\times 2$ covariance matrix
$\Gamma_{\sigma}^{(2)}=\begin{bmatrix}K_{\sigma}(0)&K_{\sigma}(h_{n})\\ K_{\sigma}(h_{n})&K_{\sigma}(0)\end{bmatrix}$.
Thus, while both methods are based on local Gaussian structures, our method relies only on scalar increments and is not a strict special case of CLE.
Theorem 3.1.
Suppose the assumptions A1–A3 and B1–B5. Then the $M$-estimator $\widehat{\theta}_{n}$ is consistent to $\theta_{0}$:
$\widehat{\theta}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\theta_{0},\quad n\to\infty.$
under the sampling (2.2).
Remark 3.3.
In Theorem 3.1, we rely on the identifiability conditions B4 and B5, which are standard assumptions for establishing consistency results. However, in the case of B5, it is often difficult to identify all the parameters in the kernel function, since the contrast function exploits only the variance structure; see the examples in the next section. Therefore, some of the parameters must be estimated separately using alternative methods. Later, we describe a method of moments approach, which also yields consistent estimators with a favorable rate of convergence.
Theorem 3.2.
Suppose the same conditions as in Theorem 3.1. Suppose further B6 and B8. Then, the estimator $\widehat{\theta}_{n}$ is asymptotically normal under (2.2):
$D_{n}(\widehat{\theta}_{n}-\theta_{0})\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}\mathcal{N}(0,\Sigma(\theta_{0})),$
where $D_{n}:=\mathrm{diag}\left(h_{n}^{-1/2}I_{p},\sqrt{n}I_{q}\right)$ and the Fisher-type information matrix is
$\Sigma(\theta_{0}):=\begin{pmatrix}\displaystyle-\partial_{t}^{2}K_{\sigma_{0}}(0)\left[2\int_{0}^{\infty}\{\partial_{\xi}\mu_{\xi}(t)\}^{\otimes 2}\,\mathrm{d}t\right]^{-1}&0\\ 0&V_{2}^{-1}(\sigma_{0})V_{1}(\sigma_{0})V_{2}^{-1}(\sigma_{0})\end{pmatrix}$.
where
$V_{1}(\sigma)=\left(\frac{1}{2}\partial_{\sigma}\log\left(-\partial_{t}^{2}K_{\sigma}(0)\right)\right)^{\otimes 2},\quad V_{2}(\sigma)=\partial_{\sigma}^{2}\log\left(-\partial_{t}^{2}K_{\sigma}(0)\right).$
Remark 3.4.
The rate of convergence for the estimator $\widehat{\xi}_{n}$ is not the standard $\sqrt{nh_{n}}$, but rather the nonstandard rate $h_{n}^{-1/2}$. Since $\sqrt{nh_{n}}/h_{n}^{-1/2}=\sqrt{nh_{n}^{2}}$, this implies that the present rate is faster under the standard high-frequency condition $nh_{n}^{2}\to 0$. However, this does not mean that $\xi_{0}$ is estimable from samples over bounded intervals under the condition $h_{n}\to 0$ alone. In fact, the condition $nh_{n}\to\infty$ is essential, as the nonstandard rate essentially originates from the DRI property; see Remark 2.3. In contrast, under a framework where standard Riemann approximation applies, such as
$\frac{1}{n}\sum_{i=1}^{n}\mu_{\xi}(t_{i-1})\to\int_{0}^{1}\mu_{\xi}(s)\,\mathrm{d}s\quad(n\to\infty),$
one would expect the standard convergence rate $\sqrt{nh_{n}}$ to be recovered. This is typically the case in non-ergodic or small-noise models.
3.2 Moment estimators
Suppose that an estimator of $\xi_{0}$ is given, say $\widehat{\xi}_{n}$; see Remark 3.1, and let
$Y_{i}^{n}=X_{t_{i}}-\int_{0}^{t_{i}}\mu_{\widehat{\xi}_{n}}(s)\,\mathrm{d}s,\quad i=1,2,\dots,n.$
Consider the following $\mathbb{R}^{q}$-valued estimating functions: for $f=(f_{1},\dots,f_{q}):\mathbb{R}\to\mathbb{R}^{q}$,
$\Phi_{n}(\sigma)$ $=\frac{1}{n}\sum_{i=1}^{n}f(Y_{i-1}^{n})-\int_{\mathbb{R}}f(z)\phi_{K_{\sigma}(0)}(z)\,\mathrm{d}z.$ (3.3)
The $Z$-estimator is given by
$\Phi_{n}(\widetilde{\sigma}_{n})=0\quad(k=1,2).$ (3.4)
Since it follows by Lemmas B.5 that, for suitable functions $f$ and $G$,
$\Phi_{n}(\sigma)\stackrel{{\scriptstyle p}}{{\longrightarrow}}\Phi(\sigma)$ $:=\int_{\mathbb{R}^{q}}f(z)\left[\phi_{K_{\sigma_{0}}(0)}(z)-\phi_{K_{\sigma}(0)}(z)\right]\,\mathrm{d}z;$ (3.5)
as $n\to\infty$ uniformly in $\sigma\in\overline{\Pi}$. Then, $\widetilde{\sigma}_{n}$ can be consistent to $\sigma_{0}$ under suitable reguralities.
Theorem 3.3.
Let $f:\mathbb{R}\to\mathbb{R}^{q}$ be a measurable function such that $f\in C^{1}(\mathbb{R})$ and there exists $C>0$ such that
$|f(x)|+|\partial_{x}f(x)|\lesssim 1+|x|^{C}.$
Suppose the assumptions A1–A3 and B8 hold, and that a consistent estimator $\widehat{\xi}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\xi_{0}$ is given. Moreover, suppose the following identifiability condition is satisfied:
$\inf_{\sigma\in\overline{\Pi}:|\sigma-\sigma_{0}|>\varepsilon}|\Phi(\sigma)|>0\quad\text{for all }\varepsilon>0.$ (3.6)
Then the $Z$-estimator $\widehat{\sigma}_{n}$ defined by (3.4) is consistent:
$\widetilde{\sigma}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\sigma_{0},\quad n\to\infty.$
Theorem 3.4.
Suppose the same assumptions as in Theorem 3.3, and that the function $f:\mathbb{R}\to\mathbb{R}^{q}$ is uniformly bounded and of polynomial growth. Suppose further that the limiting function $\Phi:\mathbb{R}^{q}\to\mathbb{R}^{q}$ is continuously differentiable at $\sigma_{0}\in\Pi$, and that the Jacobian matrix $A:=\partial_{\sigma}\Phi(\sigma_{0})\in\mathbb{R}^{q\times q}$ is invertible. Moreover, suppoe that the following limit exists:
$\Gamma^{2}:=\lim_{n\to\infty}n\,\mathrm{Var}\left(\Phi_{n}(\sigma_{0})\right)\in\mathbb{R}^{q\times q}.$
Then, the asymptotic normality holds true:
$\sqrt{n}\left(\widetilde{\sigma}_{n}-\sigma_{0}\right)\xrightarrow{d}\mathcal{N}\left(\mbox{\boldmath$0$},\ A^{-1}\Gamma^{2}A^{-\top}\right),\quad n\to\infty.$
Remark 3.5 (Alternative moment-based estimators using paired observations).
Beyond the moment function $\Phi_{n}(\sigma)$ based on single-time statistics, one can construct alternative moment-type estimators using local pairs of de-trended observations $(Y_{i-1}^{n},Y_{i}^{n})$. Lemma B.6 provides a general convergence result for statistics of the form
$\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}G(Y^{n}_{i-1},Y^{n}_{i})$ $\stackrel{{\scriptstyle p}}{{\longrightarrow}}\frac{\partial^{2}K(0)}{2K(0)}\int_{\mathbb{R}}\left[\partial_{y}G(z,z)\,z-\partial_{y}^{2}G(z,z)K(0)\right]\,\phi_{K(0)}(z)\,\mathrm{d}z,$
uniformly in $\theta\in\overline{\Theta}$ under $h_{n}\to 0$ as $n\to\infty$, where $Y_{i}^{n}=X_{t_{i}}-\int_{0}^{t_{i}}\mu_{\widehat{\xi}_{n}}(s)\,\mathrm{d}s$ and $\widehat{\xi}_{n}$ is a consistent estimator for $\xi_{0}$: $\widehat{\xi}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\xi_{0}$. The function $G$ is a smooth function of polynomial growth with $G(x,x)=0$. For example:
  • $G(x,y)=(y-x)^{2}$ yields the second-order increment moment.
  • $G(x,y)=(y-x)y^{2}$ combines local variation and level.
Such functionals can be used to construct moment equations for estimating kernel parameters. Consistency follows directly from Lemma B.6; see Section 4.1.
Remark 3.6 (Estimation of the fourth derivative $\partial_{t}^{4}K(0)$).
In the moment-based approach discussed so far, only the quantities $K(0)$ and $\partial_{t}^{2}K(0)$ can be extracted directly. However, if the kernel $K$ depends on three or more parameters, as in the case of rational quadratic kernels, higher-order information such as $\partial_{t}^{4}K(0)$ becomes essential for parameter identification and estimation. Under $K\in C^{4}([0,\infty))$, the following procedure provides a general and practical way to incorporate such information.
Let us define $\delta:=-\partial_{t}^{2}K(0)>0$, so that
$\mathbb{E}[(\Delta_{i}^{n}Z)^{2}]=2\{K(0)-K(h_{n})\}=\delta h_{n}^{2}-\frac{1}{12}\partial_{t}^{4}K(0)h_{n}^{4}+o(h_{n}^{4}),\quad h_{n}\to 0.$
Noting that $\Delta_{i}^{n}Z\sim\mathcal{N}(0,\mathbb{E}[(\Delta_{i}^{n}Z)^{2}])$, we obtain by Gaussianity
$\mathbb{E}[(\Delta_{i}^{n}Z)^{4}]=3\left(\mathbb{E}[(\Delta_{i}^{n}Z)^{2}]\right)^{2}=3\delta^{2}h_{n}^{4}-\frac{1}{2}\delta\partial_{t}^{4}K(0)h_{n}^{6}+o(h_{n}^{6}).$
Hence, the fourth derivative $\partial_{t}^{4}K(0)$ is identified by
$\partial_{t}^{4}K(0)=\frac{2}{\delta h_{n}^{2}}\left\{3\delta^{2}-\mathbb{E}\left[\left(\frac{\Delta_{i}^{n}Z}{h_{n}}\right)^{4}\right]\right\}+o(1).$
Then, by Lemma B.1, the quantity $\delta=-\partial_{t}^{2}K(0)$ is consistently estimated by
$\widehat{\delta}_{n}:=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\left(\Delta_{i}^{n}X-\mu_{\widehat{\xi}_{n}}(t_{i-1})h_{n}\right)^{2}.$
Using this, we define the estimator of $\partial_{t}^{4}K(0)$ as
$\widehat{\partial_{t}^{4}K(0)}:=\frac{1}{h_{n}^{2}}\left\{3\widehat{\delta}_{n}-\frac{1}{nh_{n}^{4}\widehat{\delta}_{n}}\sum_{i=1}^{n}\left(\Delta_{i}^{n}X-\mu_{\widehat{\xi}_{n}}(t_{i-1})h_{n}\right)^{4}\right\}.$ (3.7)
Then, this estimator is consistent as $n\to\infty$, $h_{n}\to 0$, and $nh_{n}\to\infty$.
4 Examples and simulations
4.1 Drifted Gaussian processes with Gaussian Kernels
Consider a model (2.1) with
$\mu_{\xi}(s)=\xi w(s),\quad K_{\sigma}(t)=\alpha\exp\left(-\frac{\beta}{2}t^{2}\right),$
where $w:[0,\infty)\to\mathbb{R}$, is a known function, directly Riemann integrable function, and $\xi\in\mathbb{R}$, $\gamma:=\alpha\beta\in\mathbb{R}_{+}^{2}$. Note that, in this model, we can not identify $\alpha_{0}$ and $\beta_{0}$ separately, but only $\gamma_{0}:=\alpha_{0}\beta_{0}$ because $\partial^{2}_{t}K_{\sigma}(0)=-\alpha\beta$. Then, our local-Gauss contrast function is given by
$\ell_{n}(\xi,\sigma)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{\left(\Delta_{i}^{n}X-\xi w_{i-1}h_{n}\right)^{2}}{2\alpha[1-e^{-\beta h_{n}^{2}/2}]}+\log\left(2h_{n}^{-2}\alpha[1-e^{-\beta h_{n}^{2}/2}]\right)$
$=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\frac{\left(\Delta_{i}^{n}X-\xi w_{i-1}h_{n}\right)^{2}}{\gamma}+\log\gamma+o_{p}(1),$
where $w_{i-1}=w(t_{i-1})$. Hence, we obtain the following $M$-estimator by solving the estimating equation $\nabla\ell_{n}(\theta)=0$:
$\widehat{\xi}_{n}=\frac{\sum_{i=1}^{n}w_{i-1}\Delta_{i}^{n}X}{h_{n}\sum_{i=1}^{n}w_{i-1}^{2}},\quad\widehat{\gamma}_{n}=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\left(\Delta_{i}^{n}X-\widehat{\xi}_{n}w_{i-1}h_{n}\right)^{2},$ (4.1)
which are asymptotically normal estimators for $\xi_{0}$ and $\gamma_{0}$, respectively.
For separate estimation of $\alpha$ and $\beta$, we shall consider the method of moment. For example, we can use Lemmas B.5 and B.6 with $f(x,\theta)=x^{2}$ and $G(x,y,\theta)=(y-x)^{2}$ or $(y-x)y^{2}$, respectively: it follows for
$Y_{i}^{n}:=X_{t_{i}}-\widehat{\xi}_{n}\int_{0}^{t_{i}}w(s)\,\mathrm{d}s;\quad\Delta_{i}^{n}Y:=Y_{i}^{n}-Y_{i-1}^{n},$
that
$\frac{1}{n}\sum_{i=1}^{n}(Y_{i-1}^{n})^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\alpha;\quad\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}(\Delta_{i}^{n}Y)^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\alpha\beta;\quad\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\Delta_{i}^{n}Y\cdot(Y_{i-1}^{n})^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}3\alpha^{2}\beta,$
as $n\to\infty$. For example, using the first convergence, we can estimate $\alpha$ and $\beta$ separately by, for example,
$\widehat{\alpha}_{n}=\frac{1}{n}\sum_{i=1}^{n}(Y_{i-1}^{n})^{2},\quad\widehat{\beta}_{n}=\frac{\sum_{i=1}^{n}(\Delta_{i}^{n}Y)^{2}}{h_{n}^{2}\sum_{i=1}^{n}(Y_{i-1}^{n})^{2}}\ \left(=\frac{\widehat{\gamma}_{n}}{\widehat{\alpha}_{n}}\right),$ (4.2)
both of which are asymptotically normal with the rate $\sqrt{n}$.
Remark 4.1.
In this model, the local-Gauss contrast function depends only on the product $g=\alpha\beta$ through the expansion
$K_{\sigma}(0)-K_{\sigma}(h_{n})=\alpha\left(1-e^{-\beta h_{n}^{2}/2}\right)=\tfrac{1}{2}\alpha\beta h_{n}^{2}+o(h_{n}^{3}),$
as $h_{n}\to 0$. Therefore, $\alpha$ and $\beta$ are not separately identifiable from the contrast function alone. This issue is resolved by the method of moments, which utilizes higher-order statistics of the de-trended process $Y_{i}^{n}$. Alternatively, composite likelihood methods such as Bennedsen et al. [2] use multivariate Gaussian densities over $q$-dimensional blocks (e.g., $q=3$), incorporating multiple lagged covariances like $K(h_{n})$ and $K(2h_{n})$. This richer information structure enables the separate identification of $\alpha$ and $\beta$ through nonlinear relationships among the covariances.
4.2 Ornstein-Uhlenbeck processes
[…]
$K_{\sigma}(t)=\alpha e^{-\beta|t|},$
and the target parameter is $\theta:=\sigma=(\alpha,\beta)$.
As is well known, the process $X$ satisfies the following stochastic differential equation:
$\mathrm{d}X_{t}=-\beta X_{t}\,\mathrm{d}t+\sqrt{2\alpha\beta}\mathrm{d}W_{t},\quad X_{0}=Z_{0},$ (4.3)
for a Wiener process $W$. In the context of inference for SDEs, under the asymptotic regime $h_{n}\to 0$, $nh_{n}\to\infty$, and $nh_{n}^{3}\to 0$, local-Gaussian contrast by Kessler [11]:
$l_{n}(\alpha,\beta)=-\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{(\Delta_{i}^{n}X+\beta h_{n}X_{t_{i-1}})^{2}}{4\alpha\beta h_{n}}+\log(4\pi\alpha\beta h_{n})\right\},$
gives an asymptotically efficient estimator for $\alpha$ and $\beta$. If $\alpha$ is known, we can use the following asymptotically equivalent contrast for $\beta$:
$l_{n}(\alpha,\beta)=-\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{(\Delta_{i}^{n}X)^{2}}{4\alpha\beta h_{n}}+\log(4\pi\alpha\beta h_{n})\right\}.$
Now, we suppose that $\alpha$ is known. Then an asymptotic efficient estimator of $\beta$ given by the maximizer of the last contrast function: if $\alpha$ is known, then
$\widetilde{\beta}_{n}$ $=\frac{1}{4\alpha nh_{n}}\sum_{i=1}^{n}(\Delta_{i}^{n}X)^{2},\quad n\to\infty.$ (4.4)
and it holds that
$\sqrt{n}(\widetilde{\beta}_{n}-\beta)\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}\mathcal{N}(0,2\beta^{2}),$
which is an benchmark of the estimator.
We shall reconsider this estimation from the view point of Gaussian processes. First, we can use Lemma A.5 with $f(z)=z^{2}$ to estimate $\alpha$:
$\widehat{\alpha}_{n}=\frac{1}{n}\sum_{i=1}^{n}(X_{t_{i-1}})^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\alpha,$ (4.5)
which is also asymptotically normal by Theorem 3.4.
Because the O-U kernel is not smooth as in A2, we will use a mollifier $\varphi_{\epsilon}$ with $\int\varphi_{\epsilon}(s)\,\mathrm{d}s=1$, and approximate $K$ by the smoothed kernel as in Example 2.2:
$K_{\sigma}^{({\epsilon})}(t):=\int_{\mathbb{R}}K_{\sigma}(t-s)\varphi_{\epsilon}(s)\,\mathrm{d}s.$
That is, instead of modeling the data by an exact Ornstein–Uhlenbeck process, one can model it by a Gaussian process with kernel $K_{\sigma}^{(\varepsilon)}$ using a small $\varepsilon>0$.
The contrast function is
$\ell_{n}^{({\epsilon})}(\alpha,\beta)=\frac{S_{n}}{2V^{({\epsilon})}(\alpha,\beta)}+\log\left(2h_{n}^{-2}V^{({\epsilon})}(\alpha,\beta)\right).$
where $S_{n}:=\frac{1}{n}\sum_{i=1}^{n}(\Delta_{i}^{n}X)^{2}$ and $V^{({\epsilon})}(\sigma):=K_{\sigma}^{({\epsilon})}(0)-K_{\sigma}^{({\epsilon})}(h_{n})$. The score function is given by
$\frac{\partial\ell_{n}^{({\epsilon})}}{\partial\alpha}$ $=\left(1-\frac{S_{n}}{2V^{({\epsilon})}(\alpha,\beta)}\right)\cdot\frac{\partial}{\partial\alpha}\log V^{({\epsilon})}(\alpha,\beta),$
$\frac{\partial\ell_{n}^{({\epsilon})}}{\partial\beta}$ $=\left(1-\frac{S_{n}}{2V^{({\epsilon})}(\alpha,\beta)}\right)\cdot\frac{\partial}{\partial\beta}\log V^{({\epsilon})}(\alpha,\beta).$
Therefore an M-estimator is given by solving the equation
$V^{({\epsilon})}(\alpha,\beta)=\frac{1}{2}S_{n}.$
To obtain an explicit estimator we shall use the Laplace mollifier $\varphi_{\epsilon}(s)=\frac{1}{2{\epsilon}}e^{-\frac{|s|}{{\epsilon}}}$. Then, thanks to Example 2.2,
$K_{\sigma}^{({\epsilon})}(t)$ $=\frac{\alpha}{1+\beta{\epsilon}}+\left(\frac{\alpha\beta^{2}}{1+\beta{\epsilon}}-\frac{\alpha^{2}\beta}{{\epsilon}}\right)t^{2}+o(t^{3}),\quad(t\to 0)$
and, as ${\varepsilon}^{-1}>\beta$,
$V^{({\epsilon})}(\alpha,\beta)$ $=\left(\frac{\alpha^{2}\beta}{{\epsilon}}-\frac{\alpha\beta^{2}}{1+\beta{\epsilon}}\right)h_{n}^{2}+o(h_{n}^{3})=\alpha\beta\frac{h_{n}^{2}}{{\epsilon}}+o\left(\frac{h_{n}^{3}}{{\epsilon}}\right).$
For example, if $\beta$ is known then $\alpha$ is identifiable (satisfying B5), and we have
$\widehat{\beta}_{n}^{({\epsilon})}$ $=\dfrac{1}{2}S_{n}\left[\widehat{\alpha}_{n}\frac{h^{2}}{{\epsilon}}+o\left(\frac{h^{3}}{{\epsilon}}\right)\right]^{-1}=\frac{{\epsilon}}{2\widehat{\alpha}_{n}nh_{n}^{2}}\sum_{i=1}^{n}(\Delta_{i}^{n}X)^{2},\quad$ (4.6)
Hence, taking
${\epsilon}=h_{n}/2\ (\to 0),$ (4.7)
our estimator can be asymptotically efficient as in (4.4). Moreover, Theorem 3.2 leads that: for any fixed ${\epsilon}>0$,
$\sqrt{n}(\widehat{\beta}_{n}^{({\epsilon})}-\beta)\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}\mathcal{N}\left(0,2\beta^{2}+r_{\epsilon}\right),\quad n\to\infty,$
and $r_{\epsilon}=O({\epsilon})$ as ${\epsilon}\to 0$. Therefore, our mollified estimator $\widehat{\beta}_{n}^{({\epsilon})}$ is ‘approximately’ asymptotically efficient for small $\epsilon$}.
4.2.1 The Rational Quadratic kernel
Consider an example that $Z$ has the Rational Quadratic (RQ) kernel given in Example 2.1:
$K_{\sigma}(t)=\alpha\left(1+\frac{\beta^{2}}{2\gamma}t^{2}\right)^{-\gamma},\quad\sigma=(\alpha,\beta,\gamma)\in\mathbb{R}_{+}^{3},$
with parameters $\sigma^{2}>0$, $\kappa>0$, and $\alpha>0$. This kernel arises as a scale mixture of squared exponential kernels and is widely used in Gaussian process modeling for its flexibility.
The spectral density associated with this kernel has a closed-form expression:
$f(\omega)=\alpha\cdot\frac{\sqrt{2\pi}\,\Gamma(\gamma+\frac{1}{2})}{\Gamma(\gamma)}\cdot\frac{1}{(\beta^{2}\gamma)^{1/2}}\left(1+\frac{2\pi^{2}\omega^{2}}{\beta^{2}\gamma}\right)^{-(\gamma+\frac{1}{2})},$
where $\Gamma(\cdot)$ denotes the gamma function. While analytically available, this density is nonlinear in all parameters and requires nontrivial numerical treatment for Whittle-type likelihood inference.
In contrast, our method only relies on the second-order behavior of the kernel at the origin. Specifically, a simple Taylor expansion yields:
$K_{\sigma}(t)=\alpha\left\{1-\frac{\beta^{2}}{2}t^{2}+\frac{(1+\gamma)\beta^{4}}{8\gamma}t^{4}+o(t^{4})\right\},\quad t\to 0.$
so that
$\partial^{2}_{t}K_{\sigma}(0)=\alpha\beta^{2}=:\delta.$
This quantity enters directly into the contrast function and can be computed in closed form regardless of $\alpha$, allowing for efficient and robust estimation. Therefore, even for kernels with analytically known but numerically complex spectral densities, our contrast-based method offers practical advantages in terms of implementation and stability.
Actually, we can construct those estimators as follows: To estimate $\xi$ and $\delta:=\alpha\beta^{2}$, the contrast function is given by
$\ell_{n}(\xi,\delta)=\frac{1}{n}\sum_{i=1}^{n}\frac{\left(\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}\right)^{2}}{2\delta h_{n}^{2}}+\log(2\delta),$
and, by minimizing this contrast function, we obtain
$\widehat{\xi}_{n}=\arg\min_{\xi\in\overline{\Xi}}\sum_{i=1}^{n}\left(\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}\right)^{2};\quad\widehat{\delta}_{n}=\frac{1}{2nh_{n}^{2}}\sum_{i=1}^{n}\left(\Delta_{i}^{n}X-\mu_{\widehat{\xi}_{n}}(t_{i-1})h_{n}\right)^{2}.$
Hence we also obtain that, for $Y_{i}^{n}:=X_{t_{i}}-\int_{0}^{t_{i}}\mu_{\widehat{\xi}_{n}}(s)\,\mathrm{d}s$,
$\widehat{\alpha}_{n}:=\frac{1}{n}\sum_{i=1}^{n}(Y_{i-1}^{n})^{2};\quad\widehat{\beta}_{n}=\sqrt{\frac{\widehat{\delta}_{n}}{\widehat{\alpha}_{n}}}.$
To estimate $\gamma$, we need the information about $\partial_{t}^{4}K_{\sigma}$, and we may use (3.7) in Remark 3.6:
$\widehat{\partial_{t}^{4}K(0)}:=\frac{1}{h_{n}^{2}}\left\{3\widehat{\delta}_{n}^{2}-\frac{1}{nh_{n}^{4}\widehat{\delta}_{n}}\sum_{i=1}^{n}\left(\Delta_{i}^{n}X-\mu_{\widehat{\xi}_{n}}(t_{i-1})h_{n}\right)^{4}\right\}.$
Noticing that
$\partial_{t}^{4}K_{\sigma}(0)=\frac{3\alpha\beta^{4}(1+\gamma)}{\gamma},$
we have the following consistent estimator
$\widehat{\gamma}_{n}=\frac{3\widehat{\alpha}_{n}\widehat{\beta}_{n}^{4}}{\widehat{\partial_{t}^{4}K(0)}-3\widehat{\alpha}_{n}\widehat{\beta}_{n}^{4}}.$
4.3 Numerical experiments
4.3.1 Drifted Gaussian processes with Gaussian Kernels
Let us consider Example 4.1:
$\mu_{\xi}(s)=\xi e^{-s},\quad K_{\sigma}(t)=\alpha\exp\left(-\frac{\beta}{2}t^{2}\right),$
with the true values of the parameter
$(\xi_{0},\alpha_{0},\beta_{0})=(2.0,1.0,1.0).$
We compute the estimators given in (4.1) and (4.2):
$\widehat{\xi}_{n}=\frac{\sum_{i=1}^{n}e^{-{t_{i-1}}}\Delta_{i}^{n}X}{h_{n}\sum_{i=1}^{n}e^{-2{t_{i-1}}}},\quad\widehat{\alpha}_{n}=\frac{1}{n}\sum_{i=1}^{n}\left(X_{t_{i-1}}-\widehat{\xi}_{n}(1-e^{-{t_{i-1}}})\right)^{2},\quad\widehat{\beta}_{n}=\frac{\widehat{\gamma}_{n}}{\widehat{\alpha}_{n}},$
where
$\widehat{\gamma}_{n}=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\left(\Delta_{i}^{n}X-\widehat{\xi}_{n}e^{-{t_{i-1}}}h_{n}\right)^{2}.$
We shall try the following two cases:
  • (I)
    $h_{n}=n^{-0.4}$, where $T_{n}:=nh_{n}=n^{0.6}\to\infty$, and in estimating $\widehat{\alpha}_{n}$ or $\widehat{\beta}_{n}$, the other parameter and $\xi$ were assumed to be known and set to their true values.
  • (II)
    The same setting as in Case (I), and all the parameters are estimated jointly (we will use $\widehat{\xi}_{n}$ in estimating $\widehat{\alpha}_{n}$ and $\widehat{\beta}_{n}$).
  • (III)
    $h_{n}=n^{-0.8}$, where $T_{n}:=n^{0.2}\to\infty$, the terminal is smaller than that of (I) and (II). Morever, in estimating $\widehat{\alpha}_{n}$ or $\widehat{\beta}_{n}$, the other parameter and $\xi$ were assumed to be known and set to their true values.
For each n=100, 1000, 3000n=100,\ 1000,\ 3000, the experiments are itterated $500$ times, and we shall show the mean and standard deviation (s.d.) for each $\widehat{\alpha}_{n},\ \widehat{\beta}_{n}$ and $\widehat{\xi}_{n}$ in Tables 1 and 2, and normal QQ-plots for each estimators in Figures 1 and 2, respectively.
We would like to compare (I) vs. (II), and (I) vs. (III).
The result of Case (I)
$n$ $\widehat{\xi}_{n}$ $\widehat{\alpha}_{n}$ $\widehat{\beta}_{n}$ 500 1.9057 1.0294 1.0210 (1.1316) (0.3167) (0.2569) 1000 1.9059 1.0093 1.0020 (1.1489) (0.2405) (0.2029) 3000 1.9279 0.9974 1.0106 (1.2021) (0.1620) (0.1459) True 2.0 1.0 1.0
Table 1: Case (I): Means and standard deviations (in parentheses) of the estimators $\widehat{\xi}_{n}$, $\widehat{\alpha}_{n}$, and $\widehat{\beta}_{n}$ over 500 replications, with $h_{n}=n^{-0.4}$ and other parameters fixed at their true values. The results illustrate good finite-sample accuracy and agreement with the asymptotic normality predicted by theory.
Figure 1: Normal QQ plots for Case (I): Scaled estimators $h_{n}^{-1/2}(\widehat{\xi}_{n}-\xi_{0})$ and $n^{1/2}(\widehat{\alpha}_{n}-\alpha_{0})$, $n^{1/2}(\widehat{\beta}_{n}-\beta_{0})$ over 500 replications. The plots show good agreement with the theoretical normal distribution.
The result of Case (II)
$n$ $\widehat{\xi}_{n}$ $\widehat{\alpha}_{n}$ $\widehat{\beta}_{n}$ 500 1.8199 2.2634 0.6274 (1.1414) (1.6769) (0.3382) 1000 1.9404 2.1516 0.6302 (1.1125) (1.5830) (0.3013) 3000 1.9476 2.1510 0.6396 (1.0970) (1.6094) (0.3028) True 2.0 1.0 1.0
Table 2: Case (II): Means and standard deviations (in parentheses) of the estimators $\widehat{\xi}_{n}$, $\widehat{\alpha}_{n}$, and $\widehat{\beta}_{n}$ over 500 replications, with $h_{n}=n^{-0.4}$ and all parameters estimated jointly. The results show noticeable upward bias in $\widehat{\alpha}_{n}$ and downward bias in $\widehat{\beta}_{n}$ due to error propagation from $\widehat{\xi}_{n}$.
Figure 2: Normal QQ plots for Case (II): Scaled estimators from joint estimation with $h_{n}=n^{-0.4}$. Upward bias in $\widehat{\alpha}_{n}$ and downward bias in $\widehat{\beta}_{n}$ are accompanied by departures from normality, especially for $\widehat{\alpha}_{n}$ although $\widehat{\xi}_{n}$ still seems to be asymptotically normal.
The result of Case (III)
$n$ $\widehat{\xi}_{n}$ $\widehat{\alpha}_{n}$ $\widehat{\beta}_{n}$ 500 1.9828 1.0236 1.0645 (1.1970) (0.8996) (0.8531) 1000 2.0446 0.9935 1.0318 (1.1977) (0.8225) (0.7393) 3000 2.0449 1.0202 1.0205 (1.1730) (0.7799) (0.6870) True 2.0 1.0 1.0
Table 3: Case (III): Means and standard deviations (in parentheses) of the estimators $\widehat{\alpha}_{n}$ and $\widehat{\beta}_{n}$ over 500 replications, with $\xi$ fixed at its true value, $h_{n}=n^{-0.8}$, and other parameters known. Consistency is improved compared to Case (II), but the slow growth of $T_{n}$ leads to noticeable deviations from normality in finite samples; see Figure 3, below.
Figure 3: Normal QQ plots for Case (III): Scaled estimators with $\xi$ fixed at its true value and $h_{n}=n^{-0.8}$. While bias is minimal, the slow growth of $T_{n}:=nh_{n}=n^{0.2}$ results in clear deviations from normality compared to Case (I), where $T_{n}=n^{0.6}$, reflecting insufficient mixing in finite samples.
4.3.2 Discussion of numerical experiments
In this section, we examined the finite-sample performance of the proposed estimators through simulation experiments. Theoretically, the convergence rate of the estimator for $\xi$ is $h_{n}^{-1/2}$, while the convergence rates of the estimators for $\alpha$ and $\beta$ are $n^{1/2}$, and these rates do not directly depend on the observation horizon $T_{n}=nh_{n}$. However, in finite samples, the speed at which $\int_{0}^{T_{n}}\mu_{\xi}(s)\,ds$ approaches the total mass $\int_{0}^{\infty}\mu_{\xi}(s)\,ds$ affects the estimation accuracy, so the choice of $T_{n}$ is practically important. In particular, when $\mu_{\xi}$ is directly Riemann integrable (DRI), this approximation error depends on the growth rate of $T_{n}$, and when $T_{n}$ is small, noticeable bias and distributional distortion can occur in finite samples.
In Case (I) (Tables 1, Figures 1), the other parameters were treated as known, so the sample means of the estimators were close to the true values, and the asymptotic normality predicted by theory was also observed in finite samples. In contrast, in Case (II) (Tables 2, Figures 2), all parameters were estimated jointly, and the estimation error of $\widehat{\xi}_{n}$ directly affected $\widehat{\alpha}_{n}$ and $\widehat{\beta}_{n}$, resulting in pronounced finite-sample instability. In particular, $\widehat{\alpha}_{n}$ is constructed as the mean of squared residuals obtained by subtracting the drift estimate from the observed values, so the squared estimation error $(\widehat{\xi}_{n}-\xi_{0})^{2}$ enters as a positive term, leading to an upward bias in $\widehat{\alpha}_{n}$ and a downward bias in $\widehat{\beta}_{n}$. This structural bias decreases as $n$ increases, but remains non-negligible in finite samples.
In Case (III) (Tables 3, Figures 3), $\xi$ was fixed at its true value, so such bias was not observed and consistency was improved. However, with $h_{n}=n^{-0.8}$, the growth of $T_{n}=n^{0.2}$ was extremely slow, resulting in a small effective sample size. Consequently, the mixing effect required for asymptotic normality did not sufficiently operate in finite samples, and the QQ plots showed marked distributional distortion. This indicates that even when the formal convergence rates are expressed in terms of $h_{n}$ and $n$, the condition $T_{n}\to\infty$ plays an essential role in the convergence of the tail term under DRI and in the validity of the weak-dependence CLT.
From these results, it is confirmed that the stability of the estimators and the accuracy of the normal approximation in finite samples depend on both the error propagation structure and the growth rate of $T_{n}$, and that this effect is particularly pronounced in joint estimation with real data.
Future work. Possible directions for improvement are as follows. (1) Random subsample averaging (Jackknife-after-bootstrap type): Generate random subsamples to estimate $\xi$, and use the average of these estimates. This can weaken the correlation between the estimation errors of $\xi$ and $(\alpha,\beta)$ without significantly increasing the variance of $\xi$, thus maintaining stability in finite samples more effectively than simple sample splitting. (2) Orthogonalized estimating functions using lagged covariances: Design the estimating functions for $\alpha$ and $\beta$ so that their gradient with respect to $\xi$ is zero (Neyman orthogonal), thereby ensuring that the estimation error of $\xi$ does not appear as a first-order term. In particular, use covariances at sufficiently large lags, rather than short differences, to extract information on $\alpha$ while attenuating the influence of the drift. (3) One-step stabilization: Starting from stable initial values of Case (I) type (consistent at rate $h_{n}^{-1/2}$ for $\xi$ and $n^{1/2}$ for $\alpha$ and $\beta$), apply a single Newton update to the joint estimating equations. This can correct the first-order dependence between $\xi$ and $(\alpha,\beta)$ and potentially mitigate the bias imbalance observed in finite samples. The goal here is not to improve efficiency, but to achieve better error propagation alignment and numerical stability.
5 Concluding remarks
In this paper, we have proposed a contrast-based estimation framework for Gaussian processes with time-inhomogeneous drifts, observed at high frequency. The key idea is to construct a local pseudo-likelihood using scalar variances of adjacent increments, thereby avoiding any inversion of covariance matrices. The resulting estimators are simple, computationally efficient, and theoretically tractable under general ergodicity conditions.
The main contribution of this work lies in the balance between generality and feasibility. Our method covers a broad class of stationary Gaussian processes with parametric kernel functions—including Gaussian, Matérn, and rational quadratic kernels—without requiring Markovianity or spectral representations. Even when the kernel is not smooth at the origin, such as in the Ornstein–Uhlenbeck (OU) process, we can apply mollifier techniques to restore differentiability and retain asymptotic efficiency. Furthermore, when the contrast function fails to identify all kernel parameters, moment-based corrections allow us to recover identifiability without sacrificing tractability.
In particular, as discussed in Remark 3.6, higher-order derivatives of the kernel function, such as $\partial_{t}^{4}K(0)$, can be estimated consistently using the fourth empirical moment of the residuals. This extension of the moment method enables identifiability even when the kernel depends on three or more parameters. However, it should be noted that higher-order moments often lead to estimators with large variance and numerical instability. In practice, these estimators are best used as initial values for one-step estimators or other refinement procedures, rather than as final estimates themselves.
Compared to other likelihood-based approaches, our method has distinct advantages in both scope and implementation. In particular, Whittle-type methods, which operate in the frequency domain, assume a parametric model for the spectral density and are well suited for smooth, stationary processes observed over long time spans. However, they typically require pre-removal of time-dependent mean functions and assume equispaced data without missing observations. In contrast, our method accommodates time-inhomogeneous drifts directly, works under dense (but possibly irregular) sampling schemes, and requires only minimal model assumptions.
It is worth noting that our framework does not cover nonstationary Gaussian processes such as fractional Brownian motion (fBM), whose kernel functions are not integrable and whose spectral densities are often singular. Nevertheless, our method remains practically useful because many real-world applications—such as mortality forecasting or environmental time series—are well described by stationary Gaussian noise plus a deterministic trend. Moreover, inference for long-memory processes like fBM under high-frequency designs is known to be theoretically and computationally challenging.
Future work. Several directions remain for further investigation. One promising avenue is to develop a hybrid estimation procedure that combines our time-domain contrast approach with frequency-domain techniques. Specifically, one may first estimate the drift parameter via least squares using local increments, and then apply Whittle-type likelihood methods to the de-trended residuals. This two-step procedure leverages the strengths of both domains: our method efficiently handles the drift component under high-frequency sampling, while Whittle’s method can exploit the global structure of the residual process when the underlying spectral density is sufficiently smooth. A careful analysis of the impact of the first-stage estimation error on the frequency-domain inference remains an important topic for future research.
Another potential extension is to design bias-reduction strategies for slowly decaying drift functions, possibly by incorporating parametric extrapolation or higher-order correction terms into the contrast function. Furthermore, while our moment-based extensions allow identification of additional kernel parameters, a systematic study of their finite-sample performance and robustness under model misspecification would be valuable. Finally, an important open problem is to extend the framework to certain classes of nonstationary or long-memory Gaussian processes, while maintaining computational tractability and theoretical guarantees.
6 Proofs of main theorems
6.1 Proof of Theorem 3.1
First, we shall show that the contrast function (3.1) converges to a deterministic limit uniformly in $\theta\in\overline{\Theta}$.
By Lemma B.1, we obtain
$\frac{1}{n}\sum_{i=1}^{n}\frac{(\Delta_{i}^{n}X-h_{n}\mu_{\xi}(t_{i-1}))^{2}}{2h_{n}^{2}}\stackrel{{\scriptstyle p}}{{\longrightarrow}}-\partial_{t}^{2}K(0),$
uniformly in $\xi$. On the other hand, the condition B1 implies
$K_{\sigma}(0)-K_{\sigma}(h_{n})=-\frac{1}{2}\partial_{t}^{2}K_{\sigma}(0)h_{n}^{2}+o(h_{n}^{3}),$
uniformly in $\sigma$. Therefore,
$\frac{1}{2[K_{\sigma}(0)-K_{\sigma}(h_{n})]}=\frac{1}{h_{n}^{2}\partial_{t}^{2}K_{\sigma}(0)}+o(h_{n}^{-1}),$
and
$\log(h_{n}^{-2}[K_{\sigma}(0)-K_{\sigma}(h_{n})])=\log\left(-\frac{1}{2}\partial_{t}^{2}K_{\sigma}(0)\right)+o(h_{n}),$
uniformly in $\sigma$. Combining all, we obtain
$\ell_{n}(\xi,\sigma)=\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{(\Delta_{i}^{n}X-h_{n}\mu_{\xi}(t_{i-1}))^{2}}{-2h_{n}^{2}\partial_{t}^{2}K_{\sigma}(0)}+\log[-\partial_{t}^{2}K_{\sigma}(0)]\right\}+o_{p}(1),$
uniformly in $\theta$. Since the leading term converges in probability to
$\ell(\sigma):=\frac{\partial_{t}^{2}K_{\sigma_{0}}(0)}{\partial_{t}^{2}K_{\sigma}(0)}+\log[-\partial_{t}^{2}K_{\sigma}(0)].$
Hence it follows that
$\sup_{\theta\in\overline{\Theta}}\left|\ell_{n}(\xi,\sigma)-\ell(\sigma)\right|\stackrel{{\scriptstyle p}}{{\longrightarrow}}0,\quad n\to\infty.$
Second, note that $\ell(\sigma)$ is minimized if and only if $\partial_{t}^{2}K_{\sigma_{0}}(0)=\partial_{t}^{2}K_{\sigma}(0)$, which implies that $\sigma=\sigma_{0}$ by B5, that is, it follows that
$\inf_{|\sigma-\sigma_{0}|>{\epsilon}}|\ell(\sigma)|>\ell(\sigma_{0}).$
Hence, by Theorem 5.7 by van der Vaart [20], the following consistency holds true:
$\widehat{\sigma}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\sigma_{0},\quad n\to\infty.$
Next, note that
$L_{n}(\xi)=nh_{n}\left\{\ell_{n}(\xi,\widehat{\sigma}_{n})-\ell_{n}(\xi_{0},\widehat{\sigma}_{n})\right\}.$
Then we write:
$L_{n}(\xi)$ $=h_{n}\sum_{i=1}^{n}\left\{\frac{h_{n}^{2}\left\{\mu_{\xi_{0}}^{2}(t_{i-1})-\mu_{\xi}^{2}(t_{i-1})\right\}}{2[K_{\widehat{\sigma}_{n}}(0)-K_{\widehat{\sigma}_{n}}(h_{n})]}+\frac{2h_{n}(\mu_{\xi}(t_{i-1})-\mu_{\xi_{0}}(t_{i-1}))\Delta_{i}^{n}X}{2[K_{\widehat{\sigma}_{n}}(0)-K_{\widehat{\sigma}_{n}}(h_{n})]}\right\}.$
By B1 and $\widehat{\sigma}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\sigma_{0}$,
$[K_{\widehat{\sigma}_{n}}(0)-K_{\widehat{\sigma}_{n}}(h_{n})]=-\frac{1}{2}\partial_{t}^{2}K(0)h_{n}^{2}+o_{p}(h_{n}^{3}),$
so the reciprocal is
$\frac{1}{[K_{\widehat{\sigma}_{n}}(0)-K_{\widehat{\sigma}_{n}}(h_{n})]}=-\frac{2}{\partial_{t}^{2}K(0)h_{n}^{2}}+o_{p}(h_{n}^{-1}).$
Substituting, we obtain
$L_{n}(\xi)$ $=-\frac{1}{\partial_{t}^{2}K(0)}\sum_{i=1}^{n}\left\{\mu_{\xi_{0}}^{2}(t_{i-1})-\mu_{\xi}^{2}(t_{i-1})\right\}h_{n}$
$\quad-\frac{2}{\partial_{t}^{2}K(0)}\sum_{i=1}^{n}[\mu_{\xi}(t_{i-1})-\mu_{\xi_{0}}(t_{i-1})]\Delta_{i}^{n}X+o_{p}(h_{n}).$
Now decompose $\Delta_{i}^{n}X=\Delta_{i}^{n}Z+h_{n}\mu_{\xi_{0}}(t_{i-1})+r_{i}^{n}$, where $r_{i}^{n}:=\int_{t_{i-1}}^{t_{i}}(\mu_{\xi_{0}}(s)-\mu_{\xi_{0}}(t_{i-1}))\mathrm{d}s=o(h_{n})$ by continuity. Then
$[\mu_{\xi}(t_{i-1})-\mu_{\xi_{0}}(t_{i-1})]\Delta_{i}^{n}X=[\mu_{\xi}(t_{i-1})-\mu_{\xi_{0}}(t_{i-1})]\Delta_{i}^{n}Z+h_{n}[\mu_{\xi}(t_{i-1})-\mu_{\xi_{0}}(t_{i-1})]\mu_{\xi_{0}}+o(h_{n}).$
It follows from Lemma A.1 that
$\sum_{i=1}^{n}[\mu_{\xi}(t_{i-1})-\mu_{\xi_{0}}(t_{i-1})]\Delta_{i}^{n}Z\stackrel{{\scriptstyle p}}{{\longrightarrow}}0.$
Moreover, by the direct Riemann integrability B3 (so $\mu_{\xi}$ is bounded; Remark 2.2), we have that
$\sum_{i=1}^{n}[\mu_{\xi}(t_{i-1})-\mu_{\xi_{0}}(t_{i-1})]\mu_{\xi_{0}}(t_{i-1})h_{n}\to\int_{0}^{\infty}(\mu_{\xi}(s)-\mu_{\xi_{0}}(s))\mu_{\xi_{0}}(s)\,\mathrm{d}s.$
Therefore,
$L_{n}(\xi)\to\frac{1}{\partial_{t}^{2}K(0)}\int_{0}^{\infty}(\mu_{\xi}(t)-\mu_{\xi_{0}}(t))^{2}\mathrm{d}t=:L(\xi).$
By B4, $L(\xi)=0$ if and only if $\xi=\xi_{0}$. Hence $L(\xi)>0$ for all $\xi\neq\xi_{0}$. Finally, by Theorem 5.7 by van der Vaart [20] again, it follows that
$\widehat{\xi}_{n}:=\arg\min_{\xi\in\overline{\Xi}}\ell_{n}(\xi,\widehat{\sigma}_{n})\stackrel{{\scriptstyle p}}{{\longrightarrow}}\xi_{0}.$
This completes the proof. ∎
6.2 Proof of Theorem 3.2
Since $\widehat{\theta}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\theta_{0}\in\mathring{\Theta}$, we can assume, in the standard argument for asymptotic normality, that $\widehat{\theta}_{n}\in\mathring{\Theta}$ for $n$ large enough. without loss of generality.
Let $D_{n}$ be the block-diagonal scaling matrix:
$D_{n}:=\begin{pmatrix}h_{n}^{-1/2}I_{p}&0\\ 0&\sqrt{n}I_{q}\end{pmatrix}$.
Applying Taylor’s formula around $\theta_{0}$, we write
$0=\partial_{\theta}\ell_{n}(\widehat{\theta}_{n})=\partial_{\theta}\ell_{n}(\theta_{0})+\int_{0}^{1}\partial_{\theta}^{2}\ell_{n}(\theta_{n}^{*}(u))\,\mathrm{d}u\cdot(\widehat{\theta}_{n}-\theta_{0}),$
where $\theta_{n}^{*}(u):=u\widehat{\theta}_{n}+(1-u)\theta_{0}$ for some $u\in(0,1)$. Multiplying both sides by $D_{n}$, we obtain
$0=D_{n}\partial_{\theta}\ell_{n}(\theta_{0})+D_{n}\int_{0}^{1}\partial_{\theta}^{2}\ell_{n}(\theta_{n}^{*}(u))\,\mathrm{d}u\cdot(\widehat{\theta}_{n}-\theta_{0}).$
Rewriting, we get
$D_{n}(\widehat{\theta}_{n}-\theta_{0})=-\left\{C_{n}D_{n}\partial_{\theta}^{2}\ell_{n}(\theta_{n}^{*})D_{n}^{-1}\right\}^{-1}C_{n}D_{n}\partial_{\theta}\ell_{n}(\theta_{0}),$
where
$C_{n}:=\begin{pmatrix}nh_{n}I_{p}&0\\ 0&I_{q}\end{pmatrix}$.
From Lemma B.2, we have
$C_{n}D_{n}\partial_{\theta}\ell_{n}(\theta_{0})\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}\mathcal{N}(0,J(\theta_{0})),\quad J(\theta_{0}):=\begin{pmatrix}\displaystyle\frac{2\int_{0}^{\infty}\left\{\partial_{\xi}\mu_{\xi_{0}}(t)\right\}^{\otimes 2}dt}{\partial_{t}^{2}K(0)}&0\\ 0&V_{1}(\sigma_{0})\end{pmatrix}$.
It remains to show the convergence of $D_{n}\partial_{\theta}^{2}\ell_{n}(\theta_{n}^{*})D_{n}^{-1}$. Note that
$C_{n}D_{n}\partial_{\theta}^{2}\ell_{n}(\theta_{n}^{*})D_{n}^{-1}=\begin{pmatrix}nh_{n}\partial_{\xi}^{2}\ell_{n}(\theta_{n}^{*})&\sqrt{\frac{1}{nh_{n}}}\partial_{\xi}\partial_{\sigma}^{\top}\ell_{n}(\theta_{n}^{*})\\ \sqrt{nh_{n}}\partial_{\sigma}\partial_{\xi}^{\top}\ell_{n}(\theta_{n}^{*})&\partial_{\sigma}^{2}\ell_{n}(\theta_{n}^{*})\end{pmatrix}$.
and we have that
$\partial_{\xi}^{2}\ell_{n}(\theta)$ $=\frac{h_{n}^{2}}{n}\sum_{i=1}^{n}\left[\frac{\left\{\partial_{\xi}\mu_{\xi}(t_{i-1})\right\}^{\otimes 2}}{v_{n}(\sigma)}-\frac{\left\{\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}\right\}\cdot\partial_{\xi}^{2}\mu_{\xi}(t_{i-1})}{h_{n}v_{n}(\sigma)}\right];$
$\sqrt{\frac{1}{nh_{n}}}\partial_{\xi}\partial_{\sigma}\ell_{n}(\theta)$ $=\frac{1}{n\sqrt{h_{n}}}\sum_{i=1}^{n}\left[\frac{\left\{\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}\right\}\cdot\partial_{\xi}\mu_{\xi}(t_{i-1})h_{n}\cdot\partial_{\sigma}v_{n}(\sigma)}{\left\{v_{n}(\sigma)\right\}^{2}}\right];$
$\partial_{\sigma}^{2}\ell_{n}(\theta)$ $=\frac{1}{n}\sum_{i=1}^{n}\Bigg{[}\frac{\left\{\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}\right\}^{2}\cdot\left[\partial_{\sigma}^{2}\{v_{n}(\sigma)\}\cdot\{v_{n}(\sigma)\}-2\{\partial_{\sigma}v_{n}(\sigma)\}^{2}\right]}{2\left\{v_{n}(\sigma)\right\}^{3}}$
$+\frac{\partial_{\sigma}^{2}\{v_{n}(\sigma)\}}{v_{n}(\sigma)}-\frac{\left\{\partial_{\sigma}\{v_{n}(\sigma)\}\right\}^{\otimes 2}}{\left\{v_{n}(\sigma)\right\}^{2}}\Bigg{]},$
where $v_{n}(\sigma)=K_{\sigma}(0)-K_{\sigma}(h_{n})$.
As for $\partial_{\xi}^{2}\ell_{n}(\theta)$, since $\frac{h_{n}^{2}}{v_{n}(\sigma)}=\frac{2}{-\partial_{t}^{2}K_{\sigma}(0)}+o(h_{n})$ uniformly in $\sigma$ by B1, it follows that
$nh_{n}\partial_{\xi}^{2}\ell_{n}(\theta)=\left(\frac{2}{-\partial_{t}^{2}K_{\sigma}(0)}+o(h_{n})\right)\cdot h_{n}\sum_{i=1}^{n}\left[\left\{\partial_{\xi}\mu_{\xi}(t_{i-1})\right\}^{\otimes 2}-\widetilde{Y}_{i}^{n}\cdot\partial_{\xi}^{2}\mu_{\xi}(t_{i-1})\right],$
where $\widetilde{Y}_{i}^{n}:=\frac{\Delta_{i}^{n}X-h_{n}\mu_{\xi}(t_{i-1})}{h_{n}}$. Now, by the assumption B6, the first term of the summation satisfies
$h_{n}\sum_{i=1}^{n}\left\{\partial_{\xi}\mu_{\xi}(t_{i-1})\right\}^{\otimes 2}=\sum_{i=1}^{n}\left\{\partial_{\xi}\mu_{\xi}(t_{i-1})\right\}^{\otimes 2}h_{n}\to\int_{0}^{\infty}\{\partial_{\xi}\mu_{\xi}(s)\}^{\otimes 2}\mathrm{d}s,$
as $n\to\infty$ since $(\partial_{\xi}\mu_{\xi})^{2}$ is DRI. Moreover, it follows for the second term that
$h_{n}\sum_{i=1}^{n}\widetilde{Y}_{i}^{n}\cdot\partial_{\xi}^{2}\mu_{\xi}(t_{i-1})$ $=\sum_{i=1}^{n}\left[\Delta_{i}^{n}Z+\int_{t_{i-1}}^{t_{i}}\mu_{\xi_{0}}(s)\,\mathrm{d}s-h_{n}\mu_{\xi}(t_{i-1})\right]\cdot\partial_{\xi}^{2}\mu_{\xi}(t_{i-1})$
$=\sum_{i=1}^{n}\partial_{\xi}^{2}\mu_{\xi}(t_{i-1})\cdot\Delta_{i}^{n}Z+h_{n}\sum_{i=1}^{n}\frac{1}{h_{n}}\int_{t_{i-1}}^{t_{i}}\mu_{\xi_{0}}(s)\,\mathrm{d}s\cdot\partial_{\xi}^{2}\mu_{\xi}(t_{i-1})$
$\quad-\sum_{i=1}^{n}\mu_{\xi}(t_{i-1})\cdot\partial_{\xi}^{2}\mu_{\xi}(t_{i-1})h_{n}$
$\to\int_{0}^{\infty}\mu_{\xi_{0}}(s)\partial_{\xi}^{2}\mu_{\xi}(s)\mathrm{d}s-\int_{0}^{\infty}\mu_{\xi}(s)\partial_{\xi}^{2}\mu_{\xi}(s)\mathrm{d}s,$
uniformly in $\xi\in\Xi$ by Lemma A.1 and the mean value theorem under B6 and B7. Therefore,
$nh_{n}\partial_{\xi}^{2}\ell_{n}(\theta)\to\frac{2}{\partial_{t}^{2}K_{\sigma}(0)}\int_{0}^{\infty}\left[-\mu_{\xi_{0}}(s)\partial_{\xi}^{2}\mu_{\xi}(s)+\mu_{\xi}(s)\partial_{\xi}^{2}\mu_{\xi}(s)-\{\partial_{\xi}\mu_{\xi}(s)\}^{\otimes 2}\right]\mathrm{d}s,$
uniformly in $\theta\in\overline{\Theta}$. Hence we have that
$nh_{n}\partial_{\xi}^{2}\ell_{n}(\widehat{\theta}_{n})\to-\frac{2}{\partial_{t}^{2}K_{\sigma_{0}}(0)}\int_{0}^{\infty}\{\partial_{\xi}\mu_{\xi}(s)\}^{\otimes 2}\,\mathrm{d}s.$
As for $\partial_{\sigma}^{2}\ell_{n}(\theta)$, it follow by the same argument as above that
$\partial_{\sigma}^{2}\ell_{n}(\theta)$ $=\frac{1}{n}\sum_{i=1}^{n}\left[\frac{\widetilde{Y}_{i}^{n}{}^{2}\cdot\left(\partial_{\sigma}^{2}\partial_{t}^{2}K_{\sigma}(0)\cdot\partial_{t}^{2}K_{\sigma}(0)-\left(\partial_{\sigma}\partial_{t}^{2}K_{\sigma}(0)\right)^{\otimes 2}\right)}{\left(\partial_{t}^{2}K_{\sigma}(0)\right)^{3}}+o(1)\right]$
$=\left(\frac{\partial_{\sigma}^{2}\partial_{t}^{2}K_{\sigma}(0)\cdot\partial_{t}^{2}K_{\sigma}(0)-\left(\partial_{\sigma}\partial_{t}^{2}K_{\sigma}(0)\right)^{\otimes 2}}{\left(\partial_{t}^{2}K_{\sigma}(0)\right)^{3}}\right)\cdot\frac{1}{n}\sum_{i=1}^{n}\widetilde{Y}_{i}^{n}{}^{2}+o_{p}(1).$
By Lemma B.1 with $f(z,\theta)=z^{2}$, we have:
$\frac{1}{n}\sum_{i=1}^{n}\widetilde{Y}_{i}^{n}{}^{2}\to\mathbb{E}[Z^{2}]=-\partial_{t}^{2}K(0),$
and thus
$\partial_{\sigma}^{2}\ell_{n}(\theta)\to\left(\frac{\partial_{\sigma}^{2}\partial_{t}^{2}K_{\sigma}(0)\cdot\partial_{t}^{2}K_{\sigma}(0)-\left(\partial_{\sigma}\partial_{t}^{2}K_{\sigma}(0)\right)^{\otimes 2}}{\left(\partial_{t}^{2}K_{\sigma}(0)\right)^{3}}\right)(-\partial_{t}^{2}K_{\sigma_{0}}(0)).$
uniformly in $\theta\in\Theta$. Hence, we have that
$\partial_{\sigma}^{2}\ell_{n}(\widehat{\theta}_{n})$ $\stackrel{{\scriptstyle p}}{{\longrightarrow}}\frac{\partial_{\sigma}^{2}\partial_{t}^{2}K_{\sigma_{0}}(0)}{\partial_{t}^{2}K_{\sigma_{0}}(0)}-\left(\frac{\partial_{\sigma}\partial_{t}^{2}K_{\sigma_{0}}(0)}{\partial_{t}^{2}K_{\sigma_{0}}(0)}\right)^{\otimes 2}=\partial_{\sigma}^{2}\left(\log(-\partial_{t}^{2}K_{\sigma_{0}}(0))\right)=V_{2}(\sigma_{0}).$
Similarly, we see that $\frac{1}{\sqrt{nh_{n}}}\partial_{\xi}\partial_{\sigma}\ell_{n}(\theta)=o_{p}(1)$ and $\sqrt{nh_{n}}\partial_{\xi}\partial_{\sigma}\ell_{n}(\theta)=o_{p}(1)$. Hence
$D_{n}\partial_{\theta}^{2}\ell_{n}(\theta_{n}^{*})D_{n}^{-1}\stackrel{{\scriptstyle p}}{{\longrightarrow}}I(\theta_{0}):=\begin{pmatrix}\displaystyle\frac{2}{\partial_{t}^{2}K_{\sigma_{0}}(0)}\int_{0}^{\infty}\{\partial_{\xi}\mu_{\xi}(s)\}^{\otimes 2}\,\mathrm{d}s&0\\ 0&V_{2}(\sigma_{0})\end{pmatrix}$.
As a consequence, we have
$D_{n}(\widehat{\theta}_{n}-\theta_{0})\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}{\cal N}\left(\mbox{\boldmath$0$},I^{-1}(\theta_{0})J(\theta_{0})I^{-1}(\theta_{0})\right)=\mathcal{N}(0,\Sigma(\theta_{0})).$
6.3 Proof of Theorem 3.3
We may apply Lemma B.5, which yields the uniform convergence
$\sup_{\sigma\in\overline{\Pi}}\left|\Phi_{n}(\sigma)-\Phi(\sigma)\right|\stackrel{{\scriptstyle p}}{{\longrightarrow}}0.$
In addition, since $\widetilde{\sigma}_{n}$ satisfies $\Phi_{n}(\widetilde{\sigma}_{n})=0$, and since $\Phi$ satisfies the identifiability condition (3.6), it follows from standard $Z$-estimation theory:
$\widetilde{\sigma}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\sigma_{0}.$
See, e.g., Theorem 5.9 in van der Vaart [20]. ∎
6.4 Proof of Theorem 3.4
By definition, the estimator $\widetilde{\sigma}_{n}\in\mathbb{R}^{q}$ satisfies
$\Phi_{n}(\widetilde{\sigma}_{n})=\mbox{\boldmath$0$}.$
According to (the integral form of) the mean value theorem, we obtain
$\Phi_{n}(\widetilde{\sigma}_{n})-\Phi_{n}(\sigma_{0})=\left(\int_{0}^{1}\partial_{\sigma}\Phi_{n}\left(\sigma_{0}+u(\widetilde{\sigma}_{n}-\sigma_{0})\right)\mathrm{d}u\right)(\widetilde{\sigma}_{n}-\sigma_{0}).$
Hence,
$\sqrt{n}(\widetilde{\sigma}_{n}-\sigma_{0})=-\left(\int_{0}^{1}\partial_{\sigma}\Phi_{n}\left(\sigma_{0}+u(\widetilde{\sigma}_{n}-\sigma_{0})\right)\mathrm{d}u\right)^{-1}\sqrt{n}\Phi_{n}(\sigma_{0}).$
Since $\widetilde{\sigma}_{n}\to\sigma_{0}$ in probability and $\partial_{\sigma}\Phi_{n}(\sigma)$ converges uniformly in probability to $\partial_{\sigma}\Phi(\sigma)$ on a neighborhood of $\sigma_{0}$, we obtain
$\int_{0}^{1}\partial_{\sigma}\Phi_{n}\left(\sigma_{0}+u(\widetilde{\sigma}_{n}-\sigma_{0})\right)\mathrm{d}u\stackrel{{\scriptstyle p}}{{\longrightarrow}}\partial_{\sigma}\Phi(\sigma_{0})=A,$
and the inverse converges in probability to $A^{-1}$.
Next, we apply Theorem 2.1 of Neumann [16] to obtain that
$\sqrt{n}\Phi_{n}(\sigma_{0})\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}\mathcal{N}(\mbox{\boldmath$0$},\Gamma^{2}),\quad n\to\infty,$
with $\Gamma^{2}:=\lim_{n\to\infty}n\,\mathrm{Var}\left(\Phi_{n}(\sigma_{0})\right).$
Note that
$\Phi_{n}^{(j)}(\sigma_{0})=\frac{1}{n}\sum_{i=1}^{n}\left[f_{j}(Y_{i-1}^{n})-\int_{\mathbb{R}}f_{j}(z)\,\phi_{K_{\sigma_{0}}(0)}(z)\,\mathrm{d}z\right]=:\frac{1}{n}\sum_{i=1}^{n}\zeta_{i,n}^{(j)}.$
Since $Y_{i}^{n}=Z_{t_{i}}+\Delta_{i}^{n}$, where
$\Delta_{i}^{n}:=\int_{0}^{t_{i}}\left\{\mu_{\xi_{0}}(s)-\mu_{\widehat{\xi}_{n}}(s)\right\}\mathrm{d}s,$
we define the centered version
$\widetilde{\zeta}_{i,n}^{(j)}:=f_{j}(Y_{i-1}^{n})-\mathbb{E}[f_{j}(Y_{i-1}^{n})],$
so that $\mathbb{E}[\widetilde{\zeta}_{i,n}^{(j)}]=0$, and write
$\Phi_{n}^{(j)}(\sigma_{0})=\frac{1}{n}\sum_{i=1}^{n}\widetilde{\zeta}_{i,n}^{(j)}+R_{n}^{(j)},$
where
$R_{n}^{(j)}:=\frac{1}{n}\sum_{i=1}^{n}\left\{\mathbb{E}[f_{j}(Y_{i-1}^{n})]-\int_{\mathbb{R}}f_{j}(z)\phi_{K_{\sigma_{0}}(0)}(z)\mathrm{d}z\right\}.$
By a Taylor expansion and the consistency $\widehat{\xi}_{n}\to\xi_{0}$, we have $\sup_{i}|\Delta_{i}^{n}|=o_{P}(1)$ and hence
$R_{n}^{(j)}=o_{P}(n^{-1/2}).$
To apply Theorem 2.1 of Neumann [16] to $\{\widetilde{\zeta}_{i,n}^{(j)}\}_{i=1}^{n}$, we verify the following:
  • (i)
    Square integrability: There exists a constant $v_{0}>0$ such that
    $\sum_{i=1}^{n}\mathbb{E}\left[\,(\widetilde{\zeta}_{i,n}^{(j)})^{2}\,\right]\leq v_{0}\quad\text{for all }n\in\mathbb{N}.$
    This follows from the polynomial growth of $f_{j}$ and bounded moments of $Z_{t_{i}}$, together with the fact that $Y_{i}^{n}=Z_{t_{i}}+\Delta_{i}^{n}$ and $\Delta_{i}^{n}=o_{P}(1)$.
  • (ii)
    Lindeberg-type condition: For all $\varepsilon>0$,
    $\sum_{i=1}^{n}\mathbb{E}\left[\,(\widetilde{\zeta}_{i,n}^{(j)})^{2}\cdot\mathbf{1}_{\{|\widetilde{\zeta}_{i,n}^{(j)}|>\varepsilon\sqrt{n}\}}\,\right]\to 0\quad(n\to\infty).$
    This is ensured by the same growth and moment conditions as in (ii), combined with the fact that $Z_{t_{i}}$ is Gaussian.
  • (iii)
    Weak dependence (covariance inequality): There exists a sequence $\{\theta_{r}\}_{r\in\mathbb{N}}$ with $\sum_{r=1}^{\infty}\theta_{r}<\infty$ such that, for any measurable function $g:\mathbb{R}^{u}\to\mathbb{R}$ with $|g|\leq 1$,
    $\left|\mathrm{Cov}\left(g(\widetilde{\zeta}_{s_{1},n}^{(j)},\dots,\widetilde{\zeta}_{s_{u},n}^{(j)}),\widetilde{\zeta}_{t,n}^{(j)}\right)\right|\leq(\mathbb{E}[\widetilde{\zeta}_{t,n}^{(j)2}]+\mathbb{E}[\widetilde{\zeta}_{s_{u},n}^{(j)2}]+n^{-1})\cdot\theta_{r},$
    where $s_{1}<\cdots<s_{u}<s_{u}+r\leq t$.
  • (iv)
    Convergence of variance: There exists a constant $\Gamma_{j}^{2}\in(0,\infty)$ such that
    $\mathrm{Var}\left(\sum_{i=1}^{n}\widetilde{\zeta}_{i,n}^{(j)}\right)\to\Gamma_{j}^{2}.$
Verification of (i) Square integrability.
Recall that $\widetilde{\zeta}_{i,n}^{(j)}=f_{j}(Y_{i-1}^{n})-\mathbb{E}[f_{j}(Y_{i-1}^{n})]$, where $Y_{i}^{n}=Z_{t_{i}}+\delta_{i}^{n}$ with
$\delta_{i}^{n}:=\int_{0}^{t_{i}}\{\mu_{\xi_{0}}(s)-\mu_{\widehat{\xi}_{n}}(s)\}\,\mathrm{d}s.$
By Jensen’s inequality,
$\mathbb{E}\left[(\widetilde{\zeta}_{i,n}^{(j)})^{2}\right]\leq\mathbb{E}\left[f_{j}(Y_{i-1}^{n})^{2}\right].$
Under assumption (B-1), there exists $C>0$ such that $|f_{j}(x)|\lesssim 1+|x|^{C}$, hence
$f_{j}(Y_{i-1}^{n})^{2}\lesssim 1+|Z_{t_{i}}+\delta_{i}^{n}|^{2C}\lesssim 1+|Z_{t_{i}}|^{2C}+|\delta_{i}^{n}|^{2C}.$
Since $\{Z_{t_{i}}\}_{i=1}^{n}$ has uniformly bounded moments and $\sup_{i}|\delta_{i}^{n}|=o_{P}(1)$, we obtain
$\sup_{n}\sum_{i=1}^{n}\mathbb{E}\left[(\widetilde{\zeta}_{i,n}^{(j)})^{2}\right]<\infty.$
Therefore, condition (i) is satisfied.
Verification of (ii) Lindeberg-type condition.
Let $\varepsilon>0$ be arbitrary. By assumption (B-1), we have $|f_{j}(x)|\lesssim 1+|x|^{C}$ for some $C>0$, so
$|\widetilde{\zeta}_{i,n}^{(j)}|\lesssim 1+|Y_{i-1}^{n}|^{C},\quad Y_{i}^{n}=Z_{t_{i}}+\delta_{i}^{n}.$
Since $\delta_{i}^{n}=o_{P}(1)$ uniformly and $Z_{t_{i}}$ has uniformly bounded moments of all orders, we obtain for any $r>2$
$\sup_{n,i}\mathbb{E}\left[|\widetilde{\zeta}_{i,n}^{(j)}|^{r}\right]<\infty.$
Then it follows from Markov’s inequality that
$\sum_{i=1}^{n}\mathbb{E}\left[(\widetilde{\zeta}_{i,n}^{(j)})^{2}\cdot\mathbf{1}_{\{|\widetilde{\zeta}_{i,n}^{(j)}|>\varepsilon\sqrt{n}\}}\right]\lesssim n\cdot\frac{1}{n^{(r-2)/2}}\to 0\quad(n\to\infty).$
Thus, the Lindeberg-type condition (ii) is satisfied.
Verification of (iii) Weak dependence.
Fix integers $1\leq s_{1}<\dots<s_{u}<t\leq n$ and define $r:=t-s_{u}$. Let $g:\mathbb{R}^{u}\to\mathbb{R}$ be any measurable function with $|g|\leq 1$. Set
$\theta_{r}:=\frac{\displaystyle\int_{0}^{h_{n}}\int_{0}^{h_{n}}|K(rh_{n}+u-v)|\,\mathrm{d}u\,\mathrm{d}v}{\displaystyle 2\int_{0}^{h_{n}}\int_{0}^{h_{n}}|K(u-v)|\,\mathrm{d}u\,\mathrm{d}v+h_{n}}.$
We write $Y_{i}^{n}=Z_{t_{i}}+\delta_{i}^{n}$ and note that $\delta_{i}^{n}=o_{P}(1)$ uniformly in $i$. Since $f_{j}$ is of polynomial growth and differentiable, we can linearize $f_{j}(Y_{i-1}^{n})$ around $Z_{t_{i}}$, yielding
$f_{j}(Y_{i-1}^{n})=f_{j}(Z_{t_{i-1}})+R_{i}^{n},$
where the remainder $R_{i}^{n}=f_{j}(Z_{t_{i-1}}+\delta_{i}^{n})-f_{j}(Z_{t_{i-1}})$ satisfies $R_{i}^{n}=o_{P}(1)$ under the growth condition.
Let us define the auxiliary array
$X_{i}^{n}:=\frac{1}{\sqrt{n}}f_{j}(Z_{t_{i}}),\quad\text{so that }\widetilde{\zeta}_{i,n}^{(j)}=X_{i}^{n}-\mathbb{E}[X_{i}^{n}]+o_{P}(n^{-1/2}).$
By Lemma B.3 (adapted to this setting), for such triangular arrays we have
$\left|\mathrm{Cov}\left(g(X_{s_{1}}^{n},\dots,X_{s_{u}}^{n}),X_{t}^{n}\right)\right|\leq\left(\mathbb{E}[|X_{t}^{n}|^{2}]+\mathbb{E}[|X_{s_{u}}^{n}|^{2}]+\frac{1}{n}\right)\cdot\theta_{r}.$
Since the centering operation and the $o_{P}(n^{-1/2})$ term do not affect the covariance up to $o(n^{-1})$, it follows that
$\left|\mathrm{Cov}\left(g(\widetilde{\zeta}_{s_{1},n}^{(j)},\dots,\widetilde{\zeta}_{s_{u},n}^{(j)}),\widetilde{\zeta}_{t,n}^{(j)}\right)\right|\leq\left(\mathbb{E}[|\widetilde{\zeta}_{t,n}^{(j)}|^{2}]+\mathbb{E}[|\widetilde{\zeta}_{s_{u},n}^{(j)}|^{2}]+\frac{1}{n}\right)\cdot\theta_{r}+o(n^{-1}).$
Finally, if $K\in L^{1}(\mathbb{R})$, then $\sum_{r=1}^{\infty}\theta_{r}<\infty$ by Lemma B.3, and thus the weak dependence condition (iii) is satisfied.
Verification of (iv) Convergence of variance.
Recall the definition:
$\widetilde{\zeta}_{i,n}^{(j)}:=\frac{1}{\sqrt{n}}\left\{f_{j}(Y_{i-1}^{n})-\mathbb{E}[f_{j}(Y_{i-1}^{n})]\right\},\quad Y_{i}^{n}:=Z_{t_{i}}+\delta_{i}^{n},$
with $\delta_{i}^{n}:=-\int_{0}^{t_{i}}\mu_{\widehat{\xi}_{n}}(s)\,\mathrm{d}s+\int_{0}^{t_{i}}\mu_{\xi_{0}}(s)\,\mathrm{d}s=o_{P}(1)$ uniformly in $i$.
Let us define $X_{i}^{n}:=\frac{1}{\sqrt{n}}\left\{f_{j}(Z_{t_{i-1}})-\mathbb{E}[f_{j}(Z_{t_{i-1}})]\right\}$. Then we have
$\widetilde{\zeta}_{i,n}^{(j)}=X_{i}^{n}+R_{i}^{n},\quad\text{with }R_{i}^{n}:=\frac{1}{\sqrt{n}}\left\{f_{j}(Y_{i-1}^{n})-f_{j}(Z_{t_{i-1}})-\mathbb{E}[f_{j}(Y_{i-1}^{n})-f_{j}(Z_{t_{i-1}})]\right\}.$
By the smoothness and polynomial growth of $f_{j}$, and the fact that $\delta_{i}^{n}=o_{P}(1)$ uniformly, we obtain
$\sup_{1\leq i\leq n}|R_{i}^{n}|=o_{P}(n^{-1/2}),\quad\text{hence }\sum_{i=1}^{n}R_{i}^{n}=o_{P}(1).$
It follows that
$\sum_{i=1}^{n}\widetilde{\zeta}_{i,n}^{(j)}=\sum_{i=1}^{n}X_{i}^{n}+o_{P}(1),\quad\text{so that }\mathrm{Var}\left(\sum_{i=1}^{n}\widetilde{\zeta}_{i,n}^{(j)}\right)=\mathrm{Var}\left(\sum_{i=1}^{n}X_{i}^{n}\right)+o(1).$
Now consider
$\mathrm{Var}\left(\sum_{i=1}^{n}X_{i}^{n}\right)=\sum_{i=1}^{n}\mathrm{Var}(X_{i}^{n})+2\sum_{1\leq i<j\leq n}\mathrm{Cov}(X_{i}^{n},X_{j}^{n}).$
Since $X_{i}^{n}=\frac{1}{\sqrt{n}}\{f_{j}(Z_{t_{i}})-\mathbb{E}[f_{j}(Z_{t_{i}})]\}$ and $\{Z_{t_{i}}\}$ is a stationary Gaussian process, the sequence $f_{j}(Z_{t_{i}})$ is stationary and $\alpha$-mixing under appropriate conditions on $K$. Therefore, by standard results for weakly dependent stationary sequences (see e.g., Ibragimov and Rozanov (1978), Doukhan (1994)), we have the convergence
$\mathrm{Var}\left(\sum_{i=1}^{n}X_{i}^{n}\right)=\sum_{h=-(n-1)}^{n-1}\left(1-\frac{|h|}{n}\right)\mathrm{Cov}\left(f_{j}(Z_{0}),f_{j}(Z_{|h|})\right)\to\Gamma_{j}^{2},$
where
$\Gamma_{j}^{2}:=\sum_{h\in\mathbb{Z}}\mathrm{Cov}\left(f_{j}(Z_{0}),f_{j}(Z_{h})\right)\in(0,\infty),$
provided that $K\in L^{1}(\mathbb{R})$ and $f_{j}$ is of polynomial growth. Combining the above with $\sum_{i=1}^{n}R_{i}^{n}=o_{P}(1)$, we conclude
$\mathrm{Var}\left(\sum_{i=1}^{n}\widetilde{\zeta}_{i,n}^{(j)}\right)\to\Gamma_{j}^{2},$
as required. Therefore, by Theorem 2.1 of Neumann [16],
$\sqrt{n}\Phi_{n}^{(j)}(\sigma_{0})\xrightarrow{d}\mathcal{N}(0,\Gamma_{j}^{2}),$
for some variance $\Gamma_{j}^{2}$. Applying this component-wise for $j=1,\dots,q$, we obtain the vectorial convergence
$\sqrt{n}\Phi_{n}(\sigma_{0})\xrightarrow{d}\mathcal{N}(\mbox{\boldmath$0$},\Gamma^{2}).$
Combining with the convergence of $A_{n}^{-1}\stackrel{{\scriptstyle p}}{{\longrightarrow}}A^{-1}$, we conclude that
$\sqrt{n}(\widetilde{\sigma}_{n}-\sigma_{0})\xrightarrow{d}\mathcal{N}(0,A^{-1}\Gamma^{2}A^{-1}),$
as desired.
Acknowledgement. This work is partially supported by JSPS KAKENHI Grant-in-Aid for Scientific Research (C) #24K06875; Japan Science and Technology Agency CREST #JPMJCR2115.
Appendix A Limit theorems for stationary Gaussian processes
Lemma A.1.
Let $Z=(Z_{t})_{t\geq 0}\sim GP(0,K)$ with the condition A2. Then, for any function $m:[0,\infty)\to\mathbb{R}$ which is DRI, it holds that
$\sum_{i=1}^{n}m(t_{i-1})\cdot\Delta_{i}^{n}Z\stackrel{{\scriptstyle p}}{{\longrightarrow}}0,$
as $n\to\infty$.
Proof.
Let $\chi_{i}^{n}:=m(t_{i-1})\cdot\Delta_{i}^{n}Z$, and define the filtration $\mathcal{F}_{i}^{n}:=\sigma(\Delta_{1}^{n}Z,\dots,\Delta_{i}^{n}Z)$. Then $\chi_{i}^{n}$ is $\mathcal{F}_{i}^{n}$-measurable.
First, since $(\Delta_{1}^{n}Z,\dots,\Delta_{i-1}^{n}Z,\Delta_{i}^{n}Z)$ is a multivariate Gaussian vector, the conditional mean is zero:
$\mathbb{E}[\chi_{i}^{n}|\mathcal{F}_{i-1}^{n}]=m(t_{i-1})\cdot\mathbb{E}[\Delta_{i}^{n}Z|\mathcal{F}_{i-1}^{n}]=0.$
Hence,
$\sum_{i=1}^{n}\mathbb{E}[\chi_{i}^{n}|\mathcal{F}_{i-1}^{n}]=0.$
Second, since $\mathbb{E}[\chi_{i}^{n}]=0$ and noticing that $\mathrm{Var}(\Delta_{i}^{n}Z)=2[K(0)-K(h_{n})]=O(h_{n}^{2})$ by A2, it follows that
$\left\|\sum_{i=1}^{n}\mathbb{E}\left[(\chi_{i}^{n})^{2}|\mathcal{F}_{i-1}^{n}\right]\right\|_{L^{1}({\mathbb{P}})}\leq\sum_{i=1}^{n}m^{2}(t_{i-1})\cdot\mathrm{Var}(\Delta_{i}^{n}Z)=O(h_{n})\sum_{i=1}^{n}m^{2}(t_{i-1})h_{n}$
Since $m$ is DRI, which is bounded, so $m^{2}$ is DRI as well. Therefore $\sum_{i=1}^{n}m^{2}(t_{i-1})h_{n}\to\int_{0}^{\infty}m^{2}(s)\,\mathrm{d}s<\infty$. Hence, $\left\|\sum_{i=1}^{n}\mathbb{E}\left[(\chi_{i}^{n})^{2}|\mathcal{F}_{i-1}^{n}\right]\right\|_{L^{1}({\mathbb{P}})}=O(h_{n})$, which implies the convergence in probability:
$\sum_{i=1}^{n}\mathbb{E}\left[(\chi_{i}^{n})^{2}|\mathcal{F}_{i-1}^{n}\right]\stackrel{{\scriptstyle p}}{{\longrightarrow}}0.$
Then, Lemma 9 from Genon-Catalot and Jacod [6] gives the consequence.
Lemma A.2.
Let $Z=(Z_{t})_{t\geq 0}\sim GP(0,K)$ with the conditions A1 and A2. Then, there eixists the ‘mean-square derivative’ $\dot{Z}_{t}$ in the sense that
$\frac{Z_{t+h}-Z_{t}}{h}\stackrel{{\scriptstyle L^{2}}}{{\longrightarrow}}\dot{Z}_{t},\quad h\to 0,$ (A.1)
and $\dot{Z}\sim GP\left(0,-\partial^{2}_{t}K(0)\right)$. In addition, suppose A3. Then, $\dot{Z}$ is ergodic in the sense of Corollary C.1.
Proof.
See Ibragimov and Rozanov [9], Section I.7.1. ∎
Lemma A.3.
Let $Z=(Z_{t})_{t\geq 0}\sim GP(0,K)$ with the conditions A1–A3. Then, for all functions $f:\mathbb{R}\to\mathbb{R}$ which is continuous and is of polynomial growth:
$|f(x)|\leq C(1+|x|^{p}),\quad\text{for all }x\in\mathbb{R},$
for some $C>0$ and $p\geq 1$, it holds that
$\frac{1}{nh_{n}}\int_{0}^{nh_{n}}\left\{f\left(\frac{Z_{t+h_{n}}-Z_{t}}{h_{n}}\right)-f(\dot{Z}_{t})\right\}dt\stackrel{{\scriptstyle L^{1}}}{{\longrightarrow}}0,$
under $h_{n}\to 0$ and $nh_{n}\to\infty$ as $n\to\infty$.
Proof.
Since $Z$ is mean-square differentiable (A.1), also in probability, it holds that
$f\left(\frac{Z_{t+h_{n}}-Z_{t}}{h_{n}}\right)\stackrel{{\scriptstyle p}}{{\longrightarrow}}f(\dot{Z}_{t}),\quad n\to\infty,$
by the continuous mapping theorem.
To conclude $L^{1}$ convergence of the integrals, we verify uniform integrability. By the growth condition on $f$, we have
$\left|f\left(\frac{Z_{t+h_{n}}-Z_{t}}{h_{n}}\right)\right|\leq C(1+|Y_{n}(t)|^{p}),$
where $Y_{n}(t):=\frac{Z_{t+h_{n}}-Z_{t}}{h_{n}}$. We estimate
$\mathbb{E}\left[|f(Y_{n}(t))|\cdot\mathbf{1}_{\{|f(Y_{n}(t))|>K\}}\right]\leq C\mathbb{E}\left[(1+|Y_{n}(t)|^{p})\cdot\mathbf{1}_{\{|Y_{n}(t)|>\lambda\}}\right],$
for $\lambda:=\left(\frac{K}{C}-1\right)^{1/p}$. By Hölder’s inequality,
$\mathbb{E}[|Y_{n}(t)|^{p}\cdot\mathbf{1}_{\{|Y_{n}(t)|>\lambda\}}]\leq\left(\mathbb{E}[|Y_{n}(t)|^{q}]\right)^{p/q}\cdot\left(\mathbb{P}(|Y_{n}(t)|>\lambda)\right)^{1-p/q},$
for $p>0$ and $q\in(1,2)$ with $1/p+1/q=1$. Since $Y_{n}(t)\to\dot{Z}_{t}$ in $L^{2}$, the sequence $\{Y_{n}(t)\}$ is bounded in $L^{q}$ uniformly in $n$, and the tail probability decays rapidly since it has a Gaussian tail. Thus the upper bound in the last right-hand side tends to 0 as $\lambda\to\infty$, uniformly in $n$. Therefore, the family $\{f(Y_{n}(t))\}_{n}$ is uniformly integrable. Hence it follows by Vitali’s convergence theorem that
$\mathbb{E}\left|f\left(\frac{Z_{t+h_{n}}-Z_{t}}{h_{n}}\right)-f(\dot{Z}_{t})\right|\to 0,\quad n\to\infty.$
Finally, using Fubini’s theorem and dominated convergence,
$\mathbb{E}\left|\frac{1}{nh_{n}}\int_{0}^{nh_{n}}\left[f\left(\frac{Z_{t+h_{n}}-Z_{t}}{h_{n}}\right)-f(\dot{Z}_{t})\right]\mathrm{d}t\right|$
$=\frac{1}{nh_{n}}\int_{0}^{nh_{n}}\mathbb{E}\left|f\left(\frac{Z_{t+h_{n}}-Z_{t}}{h_{n}}\right)-f(\dot{Z}_{t})\right|\,\mathrm{d}t\to 0,\quad n\to\infty.$
This completes the proof. ∎
Lemma A.4.
Let $Z=(Z_{t})_{t\geq 0}\sim GP(0,K)$ with A1–A3, and let $f:\mathbb{R}\times\overline{\Theta}\to\mathbb{R}$ be continuous and of polynomial growth:
$|\partial_{\theta}^{k}f(x,\theta)|\leq C(1+|x|^{p}),\quad\text{for all }x\in\mathbb{R},\ k=0,1,$ (A.2)
for some $C>0$ and $p\geq 1$. Then, it holds that
$\frac{1}{n}\sum_{i=1}^{n}f\left(\frac{\Delta_{i}^{n}Z}{h_{n}},\theta\right)\stackrel{{\scriptstyle p}}{{\longrightarrow}}\int_{\mathbb{R}}f(z,\theta)\phi_{-\partial^{2}_{t}K(0)}(z)\,\mathrm{d}z,$
uniformly in $\theta\in\overline{\Theta}$, under $h_{n}\to 0$, $nh_{n}\to\infty$ as $n\to\infty$.
Proof.
First, we shall show the convergence for each fixed $\theta\in\overline{\Theta}$.
Accroding to the stationarity of $Z$, we see that
$Y_{i}^{n}:=\frac{\Delta_{i}^{n}Z}{h_{n}}=h_{n}^{-1}(1,-1)\begin{pmatrix}Z_{t_{i}^{n}}\\ Z_{t_{i-1}^{n}}\end{pmatrix}\sim N\left(0,2[K(0)-K(h_{n})]\right).$
Hence
$Y_{i}^{n}\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}\dot{Z}_{0}\sim{\cal N}(0,-\partial^{2}K(0)),\quad n\to\infty.$
Therefore, we shall show that, for each $\theta\in\overline{\Theta}$,
$G_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}f(Y_{i-1}^{n},\theta)-\mathbb{E}[f(\dot{Z}_{0},\theta)]\stackrel{{\scriptstyle p}}{{\longrightarrow}}0,\quad n\to\infty.$
Since $Y_{i-1}^{n}\sim N\left(0,2[K(0)-K(h_{n})]\right)$ and $\dot{Z}_{0}\sim N\left(0,-\partial^{2}_{t}K(0)\right)$, it follows by the dominated convergence theorem that
$\lim_{n\to\infty}\mathbb{E}[f(Y_{i-1}^{n},\theta)]=\mathbb{E}[f(\dot{Z}_{0},\theta)]=\int_{\mathbb{R}}f(z,\theta)\phi_{-\partial^{2}_{t}K(0)}(z)\,\mathrm{d}z.$
Define $g_{\theta}(t,h):=f\left(\frac{Z_{t+h}-Z_{t}}{h}\right)$, so that $f(Y_{i-1}^{n},\theta)=g_{\theta}(t_{i-1},h_{n})$. Then, using the notation in Lemma A.3, we have
$\left|G_{n}(\theta)\right|$ $\leq\left|\frac{1}{nh_{n}}\sum_{i=1}^{n}g_{\theta}(t_{i-1},h_{n})h_{n}-\frac{1}{nh_{n}}\int_{0}^{nh_{n}}g_{\theta}(t,h_{n})\,\mathrm{d}t\right|$
$\quad+\left|\frac{1}{nh_{n}}\int_{0}^{nh_{n}}\left[g_{\theta}(t,h_{n})-f(\dot{Z}_{t},\theta)\right]\,\mathrm{d}t\right|+\left|\frac{1}{nh_{n}}\int_{0}^{nh_{n}}f(\dot{Z}_{t},\theta)\,\mathrm{d}t-\mathbb{E}[f(\dot{Z}_{0},\theta)]\right|$
$=:G_{n}^{1}+G_{n}^{2}+G_{n}^{3}.$
Note that $G_{n}^{2}\stackrel{{\scriptstyle L^{1}}}{{\longrightarrow}}0$ by Lemma A.3, and that $G_{n}^{3}\to 0\ a.s.\ (n\to\infty)$ by the ergodicity of $\dot{Z}$. To complete the proof, we show $G_{n}^{1}\to 0\ a.s.$ as $n\to\infty$.
Note that, for fixed $h>0$, the function $t\mapsto g_{\theta}(t,h)$ is continuous in $t$, since $Z$ has continuous sample paths almost surely and $f$ is continuous. On each subinterval $t_{i-1},t_{i}]$, by the mean value theorem for integrals, there exists $\tau_{i}\in[t_{i-1},t_{i}]$ such that
$\int_{t_{i-1}}^{t_{i}}g_{\theta}(t,h_{n})\,\mathrm{d}t=g_{\theta}(\tau_{i},h_{n})h_{n}\quad a.s.$
Therefore, we can write
$G_{n}^{1}=\left|\frac{1}{nh_{n}}\sum_{i=1}^{n}\left[g_{\theta}(t_{i-1},h_{n})-g_{\theta}(\tau_{i},h_{n})\right]h_{n}\right|=\frac{1}{n}\sum_{i=1}^{n}\left|g_{\theta}(t_{i-1},h_{n})-g_{\theta}(\tau_{i},h_{n})\right|.$
Since $g(t,h)$ is uniformly continuous in $t$ on compacts, and $|\tau_{i}-t_{i-1}|\leq h_{n}\to 0$ we see that
$G_{n}^{1}\leq\sup_{|s-t|\leq h_{n}}|g_{\theta}(s,h_{n})-g_{\theta}(t,h_{n})|\to 0\quad a.s.,$
as $h_{n}\to 0$.
For the uniformity of convergence, we shall show the tightness of the random functions $G_{n}(\theta)(n=1,2,\dots)$, which is confirmed via the following tightness criterion:
$\sup_{n}\mathbb{E}\left[\sup_{\theta\in\overline{\Theta}}\left|\partial_{\theta}G_{n}(\theta)\right|\right]<\infty.$
Actually, it is easy to see by the condition (A.4) that
$\sup_{\theta\in\overline{\Theta}}\left|\partial_{\theta}G_{n}(\theta)\right|\lesssim\frac{1}{n}\sum_{i=1}^{n}(1+|Y_{i-1}^{n}|^{C})+\int_{\mathbb{R}}(1+|z|^{C})\phi_{-\partial^{2}_{t}K(0)}(z)\,\mathrm{d}z,$
which is integrable uniformly in $n\in\mathbb{N}$. Hence, the proof is completed. ∎
Corollary A.1.
Under the same assumptions as in Lemma A.4, it follows for any integer $\kappa\geq 1$ that
$\frac{1}{nh_{n}^{2\kappa}}\sum_{i=1}^{n}(\Delta_{i}^{n}Z)^{2\kappa}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\frac{(2\kappa)!}{2^{\kappa}\cdot\kappa!}\left(-\partial^{2}_{t}K(0)\right)^{\kappa},$
as $n\to\infty$.
Proof.
Take $f(x)=x^{\kappa}$ in Lemma A.4, and note that $\mathbb{E}[G^{\kappa}]=\frac{(2\kappa)!}{2^{\kappa}\cdot\kappa!}\sigma^{2\kappa}$ for $G\sim{\cal N}(0,\sigma^{2})$.
Lemma A.5.
Let $Z=(Z_{t})_{t\in\mathbb{R}}$ be a centered stationary Gaussian process with the conditions A1 and A2, and let $f(x,\theta):\mathbb{R}\times\overline{\Theta}\to\mathbb{R}$ be a measureble function such that
$\sup_{\theta\in\overline{\Theta}}|\partial_{x}^{k}\partial_{\theta}^{l}f(x,\theta)|\lesssim 1+|x|^{C},$ (A.3)
for integers $k$ and $l$ with $0\leq k+l\leq 1$. Then, under the conditions that $h_{n}\to 0$ and $nh_{n}\to\infty$ as $n\to\infty$, the following hold true:
$\sup_{\theta\in\overline{\Theta}}\left|\frac{1}{n}\sum_{i=1}^{n}f(Z_{t_{i-1}^{n}},\theta)-\int_{\mathbb{R}}f(z,\theta)\phi_{K(0)}(z)\,\mathrm{d}z\right|\stackrel{{\scriptstyle p}}{{\longrightarrow}}0,$
as $n\to\infty$.
Proof.
Note that
$\mathbb{E}\left|\frac{1}{n}\sum_{i=1}^{n}f(Z_{t_{i-1}^{n}},\theta)-\frac{1}{nh_{n}}\int_{0}^{nh_{n}}f(Z_{u},\theta)\,\mathrm{d}u\right|$
$\leq\frac{1}{nh_{n}}\sum_{i=1}^{n}\mathbb{E}\int_{t_{i-1}^{n}}^{t_{i}^{n}}|f(Z_{t_{i-1}^{n}},\theta)-f(Z_{u},\theta)|\,\mathrm{d}u$
$\leq\frac{1}{nh_{n}}\sum_{i=1}^{n}\int_{t_{i-1}^{n}}^{t_{i}^{n}}\left(\mathbb{E}|Z_{u}-Z_{t_{i-1}^{n}}|^{2}\right)^{1/2}\left(\mathbb{E}\left(\int_{0}^{1}\partial_{x}f(Z_{t_{i-1}^{n}}+v(Z_{u}-Z_{t_{i-1}^{n}}),\theta)\,\mathrm{d}v\right)^{2}\right)^{1/2}\,\mathrm{d}u.$
Since
$(Z_{t},Z_{s})^{\top}\sim{\cal N}(\mbox{\boldmath$0$},\Sigma(t,s)),\quad\Sigma(t,s)=\begin{pmatrix}K(0)&K(t-s)\\ K(t-s)&K(0)\end{pmatrix}$,
we see that
$\mathbb{E}|Z_{u}-Z_{t_{i-1}^{n}}|^{2}=2\left[K(0)-K(u-t_{i-1}^{n})\right].$
Then Corollary C.1 yields the convergence in probability for each $\theta\in\overline{\Theta}$:
$\frac{1}{n}\sum_{i=1}^{n}f(Z_{t_{i-1}^{n}},\theta)\stackrel{{\scriptstyle p}}{{\longrightarrow}}\mathbb{E}[f(Z_{0},\theta)]=\int_{\mathbb{R}}f(z,\theta)\phi_{K(0)}(z)\,\mathrm{d}z.$ (A.4)
As for the uniformity, we shall confirm that
$\sup_{n\in\mathbb{N}}\mathbb{E}\left[\sup_{\theta\in\overline{\Theta}}\left|\frac{1}{n}\sum_{i=1}^{n}\partial_{\theta}f(Z_{t_{i-1}^{n}},\theta)\right|\right]<\infty,$ (A.5)
which is easy to see by the condition (A.3).
Then (A.4) and (A.5) yield the consequence. ∎
Lemma A.6.
Let $Z=(Z_{t})_{t\geq 0}\sim GP(0,K)$ with A1 and A2, and let $G(x,y):\mathbb{R}^{2}\to\mathbb{R}$ be a function such that $G(x,\cdot)\in C^{2}(\mathbb{R})$ for each $x\in\mathbb{R}$. Suppose that, for each $x\in\mathbb{R}$ and $k=1,2,\dots,l\ (l\geq 2)$, $\partial_{y}^{k}G(x,y)$ is of polynomial growth w.r.t. $y$, $G(x,x)=0$, and that there exists a constant $M>0$ such that $|\partial_{y}^{l+1}G(x,y)|\leq M$ for all $x,y\in\mathbb{R}$. Then it holds that
$\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}G(Z_{t_{i-1}^{n}},Z_{t_{i}^{n}})$ $\stackrel{{\scriptstyle p}}{{\longrightarrow}}\frac{\partial^{2}K(0)}{2K(0)}\int_{\mathbb{R}}\left[\partial_{y}G(z,z)\,z-\partial_{y}^{2}G(z,z)K(0)\right]\,\phi_{K(0)}(z)\,\mathrm{d}z,$
under $h_{n}\to 0$ as $n\to\infty$.
Proof.
Note that $Z_{t_{i}}=Z_{t_{i-1}}+\Delta_{i}^{n}Z$ to apply Taylor’s formula of $y\mapsto G(x,y)$ around $y=x$:
$G(Z_{t_{i-1}},Z_{t_{i}})$ $=G(Z_{t_{i-1}},Z_{t_{i-1}}+\Delta_{i}^{n}Z)$
$=\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot\Delta_{i}^{n}Z+\frac{1}{2}\partial_{y}^{2}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot(\Delta_{i}^{n}Z)^{2}+R_{i},$
where the remainder $R_{i}$ is given by
$R_{i}=\frac{1}{6}\partial_{y}^{3}G(Z_{t_{i-1}},Z_{t_{i-1}}+\theta_{i}\Delta_{i}^{n}Z)\cdot(\Delta_{i}^{n}Z)^{3}\quad\text{for some }\theta_{i}\in(0,1).$
Now we examine each term in the average
$\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}G(Z_{t_{i-1}},Z_{t_{i}})=:T_{1}+T_{2}+T_{3},$
where $T_{3}=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}R_{i}$;
$T_{1}$ $:=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot\Delta_{i}^{n}Z;\quad T_{2}:=\frac{1}{2nh_{n}^{2}}\sum_{i=1}^{n}\partial_{y}^{2}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot(\Delta_{i}^{n}Z)^{2}.$
We decompose $T_{1}$ as
$T_{1}=$ $=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot\mathbb{E}[\Delta_{i}^{n}Z|Z_{t_{i-1}}]$
$+\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot\left(\Delta_{i}^{n}Z-\mathbb{E}[\Delta_{i}^{n}Z|Z_{t_{i-1}}]\right)$
$=:A_{n}+B_{n}.$
Since $\mathbb{E}[\Delta_{i}^{n}Z|Z_{t_{i-1}}=z]=\rho_{n}z$ with $\rho_{n}=\frac{K(h_{n})}{K(0)}=1-\frac{1}{2}\partial_{t}^{2}K(0)h_{n}^{2}+o(h_{n}^{3})$, we have
$\mathbb{E}[\Delta_{i}^{n}Z|Z_{t_{i-1}}]=-\frac{1}{2}\partial_{t}^{2}K(0)\cdot\frac{h_{n}^{2}}{K(0)}Z_{t_{i-1}}+o(h_{n}^{3}).$
Therefore,
$A_{n}=-\frac{\partial_{t}^{2}K(0)}{2K(0)}\cdot\frac{1}{n}\sum_{i=1}^{n}\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot Z_{t_{i-1}}+o_{P}(1).$
Since the function $f(z):=\partial_{y}G(z,z)\cdot z$ is of polynomial growth and satisfies the condition (A.3), Lemma A.5 implies that
$\frac{1}{n}\sum_{i=1}^{n}\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot Z_{t_{i-1}}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\int_{\mathbb{R}}\partial_{y}G(z,z)\cdot z\,\phi_{K(0)}(z)\,\mathrm{d}z.$
Hence,
$A_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}-\frac{\partial_{t}^{2}K(0)}{2K(0)}\int_{\mathbb{R}}\partial_{y}G(z,z)\cdot z\,\phi_{K(0)}(z)\,\mathrm{d}z.$
Next, we handle the centered term $B_{n}$. Note that
$B_{n}=\frac{1}{n}\sum_{i=1}^{n}\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot\widetilde{Z}_{i}^{n},\quad\widetilde{Z}_{i}^{n}=\frac{\Delta_{i}^{n}Z-\mathbb{E}[\Delta_{i}^{n}Z|Z_{t_{i-1}}]}{h_{n}^{2}},$
where $\mathbb{E}[\widetilde{Z}_{i}^{n}]=0$ and $\mathrm{Var}(\widetilde{Z}_{i}^{n})=O(h_{n}^{2})$. According to the Schwartz inequality, we have
$B_{n}^{2}\leq\left(\frac{1}{n}\sum_{i=1}^{n}\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}})^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}|\widetilde{Z}_{i}^{n}|^{2}\right)=O_{p}(h_{n}^{2})\stackrel{{\scriptstyle p}}{{\longrightarrow}}0,\quad n\to\infty,$
by Lemma A.5.
Thus, combining both parts, we conclude that
$T_{1}\stackrel{{\scriptstyle p}}{{\longrightarrow}}-\frac{\partial_{t}^{2}K(0)}{2K(0)}\int_{\mathbb{R}}\partial_{y}G(z,z)\cdot z\,\phi_{K(0)}(z)\,\mathrm{d}z.$
As for $T_{2}$, the same argument leads us that
$\frac{1}{2nh_{n}^{2}}\sum_{i=1}^{n}\partial_{y}^{2}G(Z_{t_{i-1}},Z_{t_{i-1}})\cdot(\Delta_{i}^{n}Z)^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\frac{\partial^{2}K(0)}{2}\int_{\mathbb{R}}\partial_{y}^{2}G(z,z)\,\phi_{K(0)}(z)\,\mathrm{d}z,$
and the details are omitted.
As for $T_{3}$, since $|\partial_{y}^{3}G(x,y)|\leq M$ and $\mathbb{E}[|\Delta_{i}^{n}Z|^{3}]=O(h_{n}^{3})$, it follows from Lemma A.4 that
$\left|\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}R_{i}\right|\leq M\frac{h_{n}}{n}\sum_{i=1}^{n}\left(\frac{\Delta_{i}^{n}Z}{h_{n}}\right)^{3}=O_{p}(h_{n})\to 0.$
As a result, we obtain that
$\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}G(Z_{t_{i-1}},Z_{t_{i}})\stackrel{{\scriptstyle p}}{{\longrightarrow}}\frac{\partial^{2}K(0)}{2K(0)}\int_{\mathbb{R}}\left[\partial_{y}G(z,z)\cdot z-\partial_{y}^{2}G(z,z)\cdot K(0)\right]\phi_{K(0)}(z)\,\mathrm{d}z.$
Appendix B Auxiliary Lemmas
In this section, we assume Conditions A1–A3 without further mention.
Lemma B.1.
Let $f:\mathbb{R}\times\overline{\Theta}\to\mathbb{R}$, be continuous and of polynomial growth uniformly in $\theta\in\overline{\Theta}$:
$|\partial^{k}_{x}\partial_{\theta}^{l}f(x,\theta)|\leq C(1+|x|^{p}),\quad x\in\mathbb{R},$ (B.1)
for integers $k$ and $l$ with $0\leq k+l\leq 1$, $C>0$ and $p\geq 1$, and suppose the assumption B2 in the model (2.1). Then, it holds that
$\frac{1}{n}\sum_{i=1}^{n}f\left(\frac{\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}}{h_{n}},\theta\right)\stackrel{{\scriptstyle p}}{{\longrightarrow}}\int_{\mathbb{R}}f(z,\theta)\phi_{-\partial^{2}_{t}K(0)}(z)\,\mathrm{d}z,$
as $n\to\infty$, uniformly in $\theta\in\overline{\Theta}$.
Proof.
First, we fix $\theta\in\Theta$, and define the scaled increments:
$\widetilde{Y}_{i}^{n}(\xi):=\frac{\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}}{h_{n}},\qquad Y_{i}^{n}:=\frac{\Delta_{i}^{n}Z}{h_{n}}.$
We have the decomposition
$\frac{1}{n}\sum_{i=1}^{n}f(\widetilde{Y}_{i}^{n}(\xi),\theta)=\frac{1}{n}\sum_{i=1}^{n}f(Y_{i-1}^{n},\theta)+\frac{1}{n}\sum_{i=1}^{n}\left(f(\widetilde{Y}_{i}^{n}(\xi),\theta)-f(Y_{i-1}^{n},\theta)\right).$
We first prove that the second term converges to zero in probability uniformly in $\theta$.
By the mean value theorem, there exists $y_{i}^{n}$ between $\widetilde{Y}_{i}^{n}(\xi)$ and $Y_{i-1}^{n}$ such that
$|f(\widetilde{Y}_{i}^{n}(\xi),\theta)-f(Y_{i-1}^{n},\theta)|=|\partial_{x}f(y_{i}^{n},\theta)||\widetilde{Y}_{i}^{n}(\xi)-Y_{i-1}^{n}|.$
Using the polynomial growth condition, we have
$|\partial_{x}f(y_{i}^{n},\theta)|\leq C(1+|y_{i}^{n}|^{p})\quad\text{for some constant }C>0.$
Next, observe that
$\widetilde{Y}_{i}^{n}(\xi)-Y_{i-1}^{n}$ $=\frac{1}{h_{n}}\left(\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}-\Delta_{i}^{n}Z\right)$
$=\frac{1}{h_{n}}\left(\int_{t_{i-1}}^{t_{i}}\mu_{\xi}(s)\,\mathrm{d}s-\mu_{\xi}(t_{i-1})h_{n}\right)$
$=\frac{1}{h_{n}}\int_{t_{i-1}}^{t_{i}}(\mu_{\xi}(s)-\mu_{\xi}(t_{i-1}))\,\mathrm{d}s.$
By the continuity of $\mu_{\xi}$ and the mean value theorem, there exists $\eta_{i}^{n}\in[t_{i-1},t_{i}]$ such that
$\mu_{\xi}(s)-\mu_{\xi}(t_{i-1})=\partial_{t}\mu_{\xi}(\eta_{i}^{n})(s-t_{i-1}),$
thus, by the condition B2,
$|\widetilde{Y}_{i}^{n}(\xi)-Y_{i-1}^{n}|$ $\leq\frac{1}{h_{n}}\int_{t_{i-1}}^{t_{i}}|\partial_{t}\mu_{\xi}(\eta_{i}^{n})||s-t_{i-1}|\,\mathrm{d}s$
$\leq\frac{\sup_{t>0,\theta\in\overline{\Theta}}|\partial_{t}\mu_{\xi}(t)|}{h_{n}}\int_{0}^{h_{n}}u\,\mathrm{d}u=O_{p}(h_{n}).$
Hence, $|\widetilde{Y}_{i}^{n}(\xi)-Y_{i-1}^{n}|=O_{p}(h_{n})$ uniformly in $i$.
Moreover, since $y_{i}^{n}$ lies between $Y_{i-1}^{n}$ and $\widetilde{Y}_{i}^{n}(\xi)$, which are Gaussian variables, it follows that $|y_{i}^{n}|$ also has a Gaussian tail. This implies that
$\mathbb{E}\left[(1+|y_{i}^{n}|^{p})|\widetilde{Y}_{i}^{n}(\xi)-Y_{i-1}^{n}|\right]\leq Ch_{n},$
uniformly in $i$ and $\theta$. Therefore,
$\frac{1}{n}\sum_{i=1}^{n}|f_{\theta}(\widetilde{Y}_{i}^{n}(\xi))-f_{\theta}(Y_{i-1}^{n})|\stackrel{{\scriptstyle p}}{{\longrightarrow}}0,$
for each $\theta\in\Theta$. Moreover, the tightness of $\frac{1}{n}\sum_{i=1}^{n}|f_{\theta}(\widetilde{Y}_{i}^{n}(\xi))-f_{\theta}(Y_{i-1}^{n})|$ is also easy to see from the condition (B.1), e.g., by the same argument as in the proof of Lemma A.4. Hence the above convergence is indeed uniform in $\theta\in\overline{\Theta}$. As a consequence, Lemma A.4 completes the proof. ∎
Lemma B.2.
Suppose that Assumptions B3, B6 and B8 hold. Then, it follows for the block diagonals $D_{n}:=\mathrm{diag}\left(h_{n}^{-1/2}I_{p},I_{q}\right)$ and $C_{n}:=\mathrm{diag}\left(nh_{n}I_{p},\sqrt{n}I_{q}\right)$ that
$C_{n}D_{n}\partial_{\theta}\ell_{n}(\theta_{0})=\begin{pmatrix}n\sqrt{h_{n}}\partial_{\xi}\ell_{n}(\theta_{0})\\ \sqrt{n}\partial_{\sigma}\ell_{n}(\theta_{0})\end{pmatrix}\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}\mathcal{N}(0,J(\theta_{0})),$
where
$J(\theta_{0}):=\begin{pmatrix}-\frac{2}{\partial_{t}^{2}K(0)}\int_{0}^{\infty}\{\partial_{\xi}\mu_{\xi}(t)\}^{\otimes 2}\,\mathrm{d}t&0\\ 0&\left(\frac{1}{2}\partial_{\sigma}\log(-\partial_{t}^{2}K_{\sigma_{0}}(0)\right)^{\otimes 2}\end{pmatrix}$.
Proof.
From the contrast function:
$\ell_{n}(\xi,\sigma):=\frac{1}{n}\sum_{i=1}^{n}\left\{\frac{(\Delta_{i}^{n}X-h_{n}\mu_{\xi}(t_{i-1}))^{2}}{2[K_{\sigma}(0)-K_{\sigma}(h_{n})]}+\log\left(2h_{n}^{-2}[K_{\sigma}(0)-K_{\sigma}(h_{n})]\right)\right\},$
the score vector is given by
$\partial_{\xi}\ell_{n}(\xi,\sigma)$ $=-\frac{h_{n}}{nv_{n}(\sigma)}\sum_{i=1}^{n}(\Delta_{i}^{n}X-h_{n}\mu_{\xi}(t_{i-1}))\cdot\partial_{\xi}\mu_{\xi}(t_{i-1}),$
$\partial_{\sigma}\ell_{n}(\xi,\sigma)$ $=\sum_{i=1}^{n}\left(-\frac{(\Delta_{i}^{n}X-h_{n}\mu_{\xi}(t_{i-1}))^{2}}{2v_{n}^{2}(\sigma)}+\frac{1}{v_{n}(\sigma)}\right)\cdot\frac{1}{n}\partial_{\sigma}v_{n}(\sigma).$
where $v_{n}(\sigma):=K_{\sigma}(0)-K_{\sigma}(h_{n})$. Noticing that $\Delta_{i}^{n}X=\Delta_{i}^{n}Z+\delta_{i}^{n}$ with
$\delta_{i}^{n}:=\int_{t_{i-1}}^{t_{i}}\mu_{\xi_{0}}(s)\,\mathrm{d}s-\mu_{\xi_{0}}(t_{i-1})h_{n}=\frac{1}{2}\partial_{t}\mu_{\xi_{0}}(t_{i-1})h_{n}^{2}+o(h_{n}^{2})=O(h_{n}^{2}),$
and
$\frac{h_{n}}{v_{n}(\sigma_{0})}=\frac{2}{h_{n}\partial_{t}^{2}K(0)}+o(1).$
since $v_{n}(\sigma_{0})=\frac{1}{2}h_{n}^{2}\partial_{t}^{2}K(0)+o(h_{n}^{3})$ by A2, we obtain that
$\partial_{\xi}\ell_{n}(\theta_{0})$ $=-\left(\frac{2}{h_{n}\partial_{t}^{2}K(0)}+o(1)\right)\Bigg{[}\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}^{n}Z\partial_{\xi}\mu_{\xi_{0}}(t_{i-1})-\frac{1}{n}\sum_{i=1}^{n}\partial_{\xi}\mu_{\xi_{0}}(t_{i-1})\cdot O(h_{n}^{2})$
$\qquad+\frac{1}{n}\sum_{i=1}^{n}\mu_{\xi_{0}}(t_{i-1})\partial_{\xi}\mu_{\xi_{0}}(t_{i-1})h_{n}\Bigg{]}$
$=-\frac{2}{nh_{n}\partial_{t}^{2}K(0)}\sum_{i=1}^{n}\Delta_{i}^{n}Z\partial_{\xi}\mu_{\xi_{0}}(t_{i-1})+O_{p}\left(\frac{1}{n}\right).$
under the DRI condition B3 and B6. Now multiply both sides by $n\sqrt{h_{n}}$:
$n\sqrt{h_{n}}\cdot\partial_{\xi}\ell_{n}(\theta_{0})$ $=-\frac{2}{\partial_{t}^{2}K(0)}\sum_{i=1}^{n}\frac{1}{\sqrt{h_{n}}}\Delta_{i}^{n}Z\partial_{\xi}\mu_{\xi_{0}}(t_{i-1})+O_{p}\left(\sqrt{h_{n}}\right).$
Next, it follows that
$\sqrt{n}\partial_{\sigma}\ell_{n}(\theta_{0})$ $=\frac{\partial_{\sigma}v_{n}(\sigma_{0})}{\sqrt{n}}\sum_{i=1}^{n}\left(\frac{1}{v_{n}(\sigma_{0})}-\frac{1}{2v^{2}_{n}(\sigma_{0})}(\Delta_{i}^{n}Z)^{2}\right).$
using $\partial_{\sigma}v_{n}(\sigma_{0})=-\frac{h_{n}^{2}}{2}\partial_{t}\partial_{\sigma}K(0)+o(h_{n}^{3})$ and the above expansion of $v_{n}$, we find:
$\frac{\partial_{\sigma}v_{n}(\sigma_{0})}{v_{n}(\sigma_{0})}=\frac{\partial_{\sigma}\partial_{t}^{2}K_{\sigma_{0}}(0)}{\partial_{t}^{2}K(0)}+o(h_{n}).$
Hence,
$\partial_{\sigma}\ell_{n}(\theta_{0})=\left(\frac{\partial_{\sigma}\partial_{t}^{2}K_{\sigma_{0}}(0)}{\partial_{t}^{2}K(0)}+o(h_{n})\right)\cdot\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{(\Delta_{i}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right).$
Since each summand is $O_{p}(1)$ by Lemma A.4, the remainder term becomes $o_{p}(h_{n})$.
As a summary, letting
$X_{i}^{n}:=\sqrt{h_{n}}\Delta_{i}^{n}Z\cdot\partial_{\xi}\mu_{\xi_{0}}(t_{i-1}),\quad Y_{i}^{n}:=\frac{1}{\sqrt{n}}\left(1-\frac{(\Delta_{i}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right),$
we have that
$n\sqrt{h_{n}}\partial_{\xi}\ell_{n}(\theta_{0})$ $=-\frac{2}{\partial_{t}^{2}K(0)}\sum_{i=1}^{n}X_{i}^{n}+o_{p}(1),$
$\sqrt{n}\partial_{\sigma}\ell_{n}(\theta_{0})$ $=\frac{\partial_{\sigma}\partial_{t}^{2}K_{\sigma_{0}}(0)}{\partial_{t}^{2}K(0)}\sum_{i=1}^{n}Y_{i}^{n}+o_{p}(1).$
We now verify the conditions of Theorem 2.1 in Neumann [16] for the triangular arrays $\{X_{i}^{n}\}$ and $\{Y_{i}^{n}\}$ defined by:
$X_{i}^{n}:=\frac{1}{\sqrt{h_{n}}}\Delta_{i}^{n}Z\cdot\partial_{\xi}\mu_{\xi_{0}}(t_{i-1}),\quad Y_{i}^{n}:=\frac{1}{\sqrt{n}}\left(1-\frac{(\Delta_{i}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right).$
Now, notice that
$D_{n}\partial_{\theta}\ell_{n}(\theta_{0})$ $=\mathrm{diag}\left(-\frac{2}{\partial_{t}^{2}K(0)}I_{p},\frac{1}{\partial_{t}^{2}K(0)}I_{q}\right)\sum_{i=1}^{n}\begin{pmatrix}X_{i}^{n}\\ Y_{i}^{n}\partial_{\sigma}\partial_{t}^{2}K_{\sigma_{0}}(0)\end{pmatrix}+o_{p}(1),$
To show the weak convergence of $S_{n}$, we apply the Cramér–Wold device. That is, for any $(a,b)^{\top}\in\mathbb{R}^{p+q}$ with $a\in\mathbb{R}^{p}$, $b\in\mathbb{R}^{q}$, we denote by
$S_{n}:=\sum_{i=1}^{n}\left(a^{\top}X_{i}^{n}+b^{\top}Y_{i}^{n}\partial_{\sigma}\partial_{t}^{2}K_{\sigma_{0}}(0)\right).$
and show the weak convergence of $S_{n}$ by applying Thereom 2.1 in Neuman [16], the CLT for trriangular arrays.
(i) Mean zero.
Since $Z$ is a centered Gaussian process, $\mathbb{E}[\Delta_{i}^{n}Z]=0$ and $\mathbb{E}[(\Delta_{i}^{n}Z)^{2}]=v_{n}(\sigma_{0})$, we have $\mathbb{E}[X_{i}^{n}]=0$ and $\mathbb{E}[Y_{i}^{n}]=0$, and $\mathbb{E}[S_{n}]=0$.
(ii) Variance convergence.
For the $X$-component:
$\sum_{i=1}^{n}\mathbb{E}[(X_{i}^{n})^{\otimes 2}]$ $=h_{n}^{-1}\sum_{i=1}^{n}\mathbb{E}[(\Delta_{i}^{n}Z)^{2}]\cdot\left\{\partial_{\xi}\mu_{\xi_{0}}(t_{i-1})\right\}^{\otimes 2}$
$=v_{n}(\sigma_{0})h_{n}^{-1}\sum_{i=1}^{n}\left\{\partial_{\xi}\mu_{\xi_{0}}(t_{i-1})\right\}^{\otimes 2}\to\frac{1}{2}\partial_{t}^{2}K(0)\int_{0}^{\infty}\left\{\partial_{\xi}\mu_{\xi_{0}}(t)\right\}^{\otimes 2}dt.$
by the condition B6. For the $Y$-component:
$\sum_{i=1}^{n}\mathbb{E}[(Y_{i}^{n})^{2}]$ $=\frac{1}{n}\sum_{i=1}^{n}\operatorname{Var}\left(1-\frac{(\Delta_{i}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{4}\operatorname{Var}\left(\frac{(\Delta_{i}^{n}Z)^{2}}{v_{n}(\sigma_{0})}\right)=\frac{1}{4}\cdot\frac{1}{n}\cdot 2n=\frac{1}{2}.$
Moreover,
$\mathbb{E}[X_{i}^{n}Y_{i}^{n}]=\frac{1}{\sqrt{nh_{n}}}\cdot\mathbb{E}\left[\Delta_{i}^{n}Z\cdot\left(1-\frac{(\Delta_{i}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right)\right]\cdot\partial_{\xi}\mu_{\xi_{0}}(t_{i-1}).$
Since $\mathbb{E}[\Delta_{i}^{n}Z]=\mathbb{E}[\Delta_{i}^{n}Z^{3}]=0$, we obtain that $\mathbb{E}[X_{i}^{n}Y_{i}^{n}]=0$. As a result, we see that
$\mathbb{E}[(S_{n})^{2}]\to a^{\top}\left(\frac{2}{\partial_{t}^{2}K(0)}\int_{0}^{\infty}\left\{\partial_{\xi}\mu_{\xi_{0}}(t)\right\}^{\otimes 2}\,\mathrm{d}t\right)a+b^{\top}\frac{1}{2}\left(\partial_{\sigma}\partial_{t}^{2}K_{\sigma_{0}}(0)\right)^{\otimes 2}b.$
(i) Lyapnov condition.
Since $X_{i}^{n}$ and $Y_{i}^{n}$ are centered Gaussian (or polynomial transformations thereof), we can compute their fourth moments explicitly:
$\mathbb{E}[|X_{i}^{n}|^{4}]=\left(\frac{1}{h_{n}}\right)^{2}\cdot 3v_{n}^{2}\cdot\|\partial_{\xi}\mu(t_{i-1})\|^{4}=O\left(h_{n}^{2}\right),$
$\mathbb{E}[|Y_{i}^{n}|^{4}]=\frac{1}{n^{2}}\cdot\mathbb{E}\left[\left(1-\frac{1}{2}\chi^{2}(1)\right)^{4}\right]=O\left(\frac{1}{n^{2}}\right).$
Therefore,
$\sum_{i=1}^{n}\mathbb{E}[|S_{n,i}|^{4}]\leq 8\sum_{i=1}^{n}\left(\mathbb{E}[|a^{\top}X_{i}^{n}|^{4}]+\mathbb{E}[|b^{\top}Y_{i}^{n}|^{4}]\right)=O\left(h_{n}^{2}+\frac{1}{n^{2}}\right)\to 0.$
(ii) Mixing covariance bounds.
We apply Lemmas B.3 and B.4 below, which establish uniform bounds for the covariance of nonlinear functionals of the triangular arrays $\{X_{i}^{n}\}$ and $\{Y_{i}^{n}\}$ under the assumptions A1 and A2. In particular, for any bounded measurable function $g$ with $\|g\|_{\infty}\leq 1$, and indices $1\leq s_{1}<\dots<s_{u}<t_{1}\leq n$, the inequalities
$\left|\mathrm{Cov}\left(g(X_{s_{1}}^{n},\dots,X_{s_{u}}^{n})X_{s_{u}}^{n},\;X_{t_{1}}^{n}\right)\right|\leq\left(\mathbb{E}|X_{s_{u}}^{n}|^{2}+\mathbb{E}|X_{t_{1}}^{n}|^{2}+\frac{1}{n}\right)\cdot\theta_{r},$
$\left|\mathrm{Cov}\left(g(Y_{s_{1}}^{n},\dots,Y_{s_{u}}^{n}),\;Y_{t_{1}}^{n}Y_{t_{2}}^{n}\right)\right|\leq\left(\mathbb{E}|Y_{t_{1}}^{n}|^{2}+\mathbb{E}|Y_{t_{2}}^{n}|^{2}+\frac{1}{n}\right)\cdot\theta_{r},\quad\text{with }r=t_{1}-s_{u},$
hold for a summable sequence $\{\theta_{r}\}\in\ell^{1}$ depending on the covariance kernel $K$. This verifies condition (2.6) of Neumann [16]. The summability of $\{\theta_{r}\}$ follows from the integrability condition $K\in L^{1}(\mathbb{R})$ imposed in A2, ensuring that the dependence decays sufficiently fast to guarantee asymptotic independence in the triangular arrays.
Remark B.1.
The sequence $\{\theta_{r}\}_{r\in\mathbb{N}}$ defined in Lemma B.2 and the lemmas below depends on $n$ through the mesh size $h_{n}$. This dependence is admissible in the framework of Neumann [16], as the central limit theorem for triangular arrays of weakly dependent variables requires only that the dependence coefficients (such as $\theta_{r}$) satisfy a uniform summability condition over $n$:
$\sup_{n\in\mathbb{N}}\sum_{r=1}^{\infty}\theta_{r}^{(n)}<\infty.$
In our case, this is ensured by the integrability condition $K\in L^{1}((0,\infty))$.
Lemma B.3.
Let $\{X_{i}^{n}\}_{1\leq i\leq n}$ be the triangular array defined by
$X_{i}^{n}:=\frac{1}{\sqrt{h_{n}}}\Delta_{i}^{n}Z\cdot\partial_{\xi}\mu_{\xi_{0}}(t_{i-1}),$
where $\Delta_{i}^{n}Z:=Z_{t_{i}}-Z_{t_{i-1}}$ and define for $r\in\mathbb{N}$ the sequence
$\theta_{r}:=\frac{\displaystyle\int_{0}^{h_{n}}\!\!\int_{0}^{h_{n}}|K(rh_{n}+u-v)|\,\mathrm{d}u\,\mathrm{d}v}{\displaystyle 2\int_{0}^{h_{n}}\!\!\int_{0}^{h_{n}}|K(u-v)|\,\mathrm{d}u\,\mathrm{d}v+h_{n}}.$
Then, for any bounded measurable function $g$ with $\|g\|_{\infty}\leq 1$, and any indices $1\leq s_{1}<\dots<s_{u}<t_{1}\leq n$ with $r:=t_{1}-s_{u}$, it holds under B8 that
$\left|\mathrm{Cov}\left(g(X_{s_{1}}^{n},\dots,X_{s_{u}}^{n})X_{s_{u}}^{n},\;X_{t_{1}}^{n}\right)\right|\leq\left(\mathbb{E}|X_{s_{u}}^{n}|^{2}+\mathbb{E}|X_{t_{1}}^{n}|^{2}+\frac{1}{n}\right)\cdot\theta_{r},$
and
$\left|\mathrm{Cov}\left(g(X_{s_{1}}^{n},\dots,X_{s_{u}}^{n}),\;X_{t_{1}}^{n}X_{t_{2}}^{n}\right)\right|\leq\left(\mathbb{E}|X_{t_{1}}^{n}|^{2}+\mathbb{E}|X_{t_{2}}^{n}|^{2}+\frac{1}{n}\right)\cdot\theta_{r},$
Moreover, $\sum_{r=1}^{\infty}\theta_{r}<\infty$.
Proof.
First, note that
$X_{i}^{n}=\frac{1}{\sqrt{h_{n}}}\cdot\partial_{\xi}\mu_{\xi_{0}}(t_{i-1})\cdot\Delta_{i}^{n}Z,$
so that
$\mathrm{Cov}\left(g(\cdots)X_{s_{u}}^{n},\;X_{t_{1}}^{n}\right)=\frac{1}{h_{n}}\cdot\partial_{\xi}\mu_{\xi_{0}}(t_{s_{u}-1})\cdot\partial_{\xi}\mu_{\xi_{0}}(t_{t_{1}-1})\cdot\mathbb{E}\left[g(\cdots)\Delta_{s_{u}}^{n}Z\cdot\Delta_{t_{1}}^{n}Z\right].$
Since $|g|\leq 1$, it follows from the Cauchy-Schwarz inequality that
$\left|\mathbb{E}\left[g(\cdots)\Delta_{s_{u}}^{n}Z\cdot\Delta_{t_{1}}^{n}Z\right]\right|\leq\mathbb{E}\left|\Delta_{s_{u}}^{n}Z\cdot\Delta_{t_{1}}^{n}Z\right|\leq\int_{0}^{h_{n}}\int_{0}^{h_{n}}|K(rh_{n}+u-v)|\,\mathrm{d}u\,\mathrm{d}v.$
Similarly, the variance terms can be bounded as
$\mathbb{E}|X_{i}^{n}|^{2}\leq\frac{C_{\mu}^{2}}{h_{n}}\cdot\int_{0}^{h_{n}}\int_{0}^{h_{n}}|K(u-v)|\,\mathrm{d}u\,\mathrm{d}v,$
where $C_{\mu}$ is a constant such that $|\gamma(t)|\leq C_{\mu}$ in the assumption B8. Therefore, the desired inequality holds with $\theta_{r}$ as defined.
To prove the summability of $\{\theta_{r}\}$, we use the change of variables $x=rh_{n}+s$ with $s\in[-h_{n},h_{n}]$, and estimate
$\sum_{r=1}^{\infty}\theta_{r}\leq C\sum_{r=1}^{\infty}\int_{0}^{h_{n}}\int_{0}^{h_{n}}|K(rh_{n}+u-v)|\,\mathrm{d}u\,\mathrm{d}v\leq C^{\prime}\int_{0}^{\infty}|K(x)|\,\mathrm{d}x<\infty,$
where $C,C^{\prime}$ are constants independent of $r$.
Moreover, since
$\left|\mathrm{Cov}\left(g(\cdots),\;X_{t_{1}}^{n}X_{t_{2}}^{n}\right)\right|$ $\leq\mathbb{E}\left|g(\cdots)\cdot X_{t_{1}}^{n}X_{t_{2}}^{n}\right|\leq\mathbb{E}\left|X_{t_{1}}^{n}X_{t_{2}}^{n}\right|$
$=\frac{1}{nh_{n}}\cdot|\partial_{\xi}\mu_{\xi_{0}}(t_{t_{1}-1})|\cdot|\partial_{\xi}\mu_{\xi_{0}}(t_{t_{2}-1})|\cdot\mathbb{E}\left|\Delta_{t_{1}}^{n}Z\cdot\Delta_{t_{2}}^{n}Z\right|.$
the similar argument leads us to the consequence. ∎
Lemma B.4.
Let $\{Y_{i}^{n}\}_{1\leq i\leq n}$ be the triangular array defined by
$Y_{i}^{n}:=\frac{1}{\sqrt{n}}\left(1-\frac{(\Delta_{i}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right),$
Define for $r\in\mathbb{N}$ the sequence
$\theta_{r}:=\frac{\displaystyle\int_{0}^{h_{n}}\!\!\int_{0}^{h_{n}}|K(rh_{n}+u-v)|\,\mathrm{d}u\,\mathrm{d}v}{\displaystyle 2\int_{0}^{h_{n}}\!\!\int_{0}^{h_{n}}|K(u-v)|\,\mathrm{d}u\,\mathrm{d}v+h_{n}}.$
Then, for any bounded measurable function $g$ with $\|g\|_{\infty}\leq 1$, and any indices $1\leq s_{1}<\dots<s_{u}<t_{1}\leq n$ with $r:=t_{1}-s_{u}$, it holds that
$\left|\mathrm{Cov}\left(g(Y_{s_{1}}^{n},\dots,Y_{s_{u}}^{n})Y_{s_{u}}^{n},\;Y_{t_{1}}^{n}\right)\right|\leq\left(\mathbb{E}|Y_{s_{u}}^{n}|^{2}+\mathbb{E}|Y_{t_{1}}^{n}|^{2}+\frac{1}{n}\right)\cdot\theta_{r},$
and
$\left|\mathrm{Cov}\left(g(Y_{s_{1}}^{n},\dots,Y_{s_{u}}^{n}),\;Y_{t_{1}}^{n}Y_{t_{2}}^{n}\right)\right|\leq\left(\mathbb{E}|Y_{t_{1}}^{n}|^{2}+\mathbb{E}|Y_{t_{2}}^{n}|^{2}+\frac{1}{n}\right)\cdot\theta_{r}.$
Moreover, $\sum_{r=1}^{\infty}\theta_{r}<\infty$.
Proof.
We first observe that $Y_{i}^{n}$ can be written as
$Y_{i}^{n}=\frac{1}{\sqrt{n}}\left(1-\frac{(\Delta_{i}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right),$
so that
$\mathbb{E}|Y_{i}^{n}|^{2}=\frac{1}{n}\cdot\mathrm{Var}\left(\frac{(\Delta_{i}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right)\leq\frac{C}{n}\cdot\mathrm{Var}\left((\Delta_{i}^{n}Z)^{2}\right),$
where the last variance is controlled by the fourth moment of $\Delta_{i}^{n}Z$. Since $Z$ is a centered stationary Gaussian process, we obtain
$\mathbb{E}[(\Delta_{i}^{n}Z)^{4}]\leq C\cdot\left(\int_{0}^{h_{n}}\!\!\int_{0}^{h_{n}}|K(u-v)|\,\mathrm{d}u\,\mathrm{d}v\right)^{2},$
and hence
$\mathbb{E}|Y_{i}^{n}|^{2}\leq\frac{C^{\prime}}{n}\cdot\int_{0}^{h_{n}}\!\!\int_{0}^{h_{n}}|K(u-v)|\,\mathrm{d}u\,\mathrm{d}v.$
This upper bound motivates the choice of denominator in $\theta_{r}$ so that the quantity $\mathbb{E}|Y_{i}^{n}|^{2}$ is uniformly absorbed. Now, for $\|g\|_{\infty}\leq 1$, we apply the Cauchy–Schwarz inequality:
$\left|\mathrm{Cov}\left(g(\cdots)Y_{s_{u}}^{n},\;Y_{t_{1}}^{n}\right)\right|\leq\mathbb{E}\left|Y_{s_{u}}^{n}Y_{t_{1}}^{n}\right|.$
Using the above representation, we have
$\mathbb{E}\left|Y_{s_{u}}^{n}Y_{t_{1}}^{n}\right|=\frac{1}{n}\cdot\mathbb{E}\left|\left(1-\frac{(\Delta_{s_{u}}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right)\left(1-\frac{(\Delta_{t_{1}}^{n}Z)^{2}}{2v_{n}(\sigma_{0})}\right)\right|.$
The dominant term arises from the covariance between $(\Delta_{s_{u}}^{n}Z)^{2}$ and $(\Delta_{t_{1}}^{n}Z)^{2}$, which is controlled by
$\mathbb{E}\left|(\Delta_{s_{u}}^{n}Z)^{2}\cdot(\Delta_{t_{1}}^{n}Z)^{2}\right|\leq C\cdot\int_{0}^{h_{n}}\!\!\int_{0}^{h_{n}}|K(rh_{n}+u-v)|\,\mathrm{d}u\,\mathrm{d}v.$
Therefore, the bound with $\theta_{r}$ holds. The same reasoning applies to
$\left|\mathrm{Cov}\left(g(\cdots),\;Y_{t_{1}}^{n}Y_{t_{2}}^{n}\right)\right|\leq\mathbb{E}\left|Y_{t_{1}}^{n}Y_{t_{2}}^{n}\right|,$
which is again controlled by the same integral involving $K(rh_{n}+u-v)$. Finally, since
$\sum_{r=1}^{\infty}\theta_{r}\leq C\int_{0}^{\infty}|K(x)|\,\mathrm{d}x<\infty.$
Hence, the lemma is proved. ∎
Lemma B.5.
Let $f(x,\theta):\mathbb{R}\times\overline{\Theta}\to\mathbb{R}$ be a measureble function such that
$\sup_{\theta\in\overline{\Theta}}|\partial_{x}^{k}\partial_{\theta}^{l}f(x,\theta)|\lesssim 1+|x|^{C},$ (B.2)
for integers $k$ and $l$ with $0\leq k+l\leq 1$. Then, it holds under the assumption B8 that
$\sup_{\theta\in\overline{\Theta}}\left|\frac{1}{n}\sum_{i=1}^{n}f(Y_{i-1}^{n},\theta)-\int_{\mathbb{R}}f(z,\theta)\phi_{K(0)}(z)\,\mathrm{d}z\right|\stackrel{{\scriptstyle p}}{{\longrightarrow}}0,$
under $h_{n}\to 0$ and $nh_{n}\to\infty$ as $n\to\infty$, where $Y_{i}^{n}=X_{t_{i}}-\int_{0}^{t_{i}}\mu_{\widehat{\xi}_{n}}(s)\,\mathrm{d}s$ and $\widehat{\xi}_{n}$ is a consistent estimator for $\xi_{0}$: $\widehat{\xi}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\xi_{0}$.
Proof.
By the mean value theorem and B8, we have
$\left|\mu_{\xi_{0}}(s)-\mu_{\widehat{\xi}_{n}}(s)\right|\leq\sup_{\xi\in\Xi}|\partial_{\xi}\mu_{\xi}(s)|\cdot|\widehat{\xi}_{n}-\xi_{0}|\leq\gamma(s)\cdot|\widehat{\xi}_{n}-\xi_{0}|,$
for each $s\geq 0$. Hence,
$\sup_{i\leq n}|Y_{i-1}^{n}-Z_{t_{i-1}}|\leq|\widehat{\xi}_{n}-\xi_{0}|\cdot\int_{0}^{t_{n}}\gamma(s)\,\mathrm{d}s\leq|\widehat{\xi}_{n}-\xi_{0}|\cdot\|\gamma\|_{L^{1}([0,\infty))}\stackrel{{\scriptstyle p}}{{\longrightarrow}}0.$
Now, using the growth condition (B.2) and the mean value theorem again, we obtain
$|f(Y_{i-1}^{n},\theta)-f(Z_{t_{i-1}},\theta)|\lesssim|Y_{i-1}^{n}-Z_{t_{i-1}}|\cdot\left(1+|Z_{t_{i-1}}|^{C}+|Y_{i-1}^{n}|^{C}\right).$
Since $Z_{t_{i-1}}$ has finite moments and $Y_{i-1}^{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}Z_{t_{i-1}}$ uniformly in $i$, it follows by dominated convergence that
$\sup_{\theta\in\overline{\Theta}}\left|\frac{1}{n}\sum_{i=1}^{n}f(Y_{i-1}^{n},\theta)-\frac{1}{n}\sum_{i=1}^{n}f(Z_{t_{i-1}},\theta)\right|\stackrel{{\scriptstyle p}}{{\longrightarrow}}0.$
Finally, Lemma A.5 implies
$\sup_{\theta\in\overline{\Theta}}\left|\frac{1}{n}\sum_{i=1}^{n}f(Z_{t_{i-1}},\theta)-\int f(z,\theta)\,\phi_{K(0)}(z)\,\mathrm{d}z\right|\stackrel{{\scriptstyle p}}{{\longrightarrow}}0,$
and the claim follows by the triangle inequality. ∎
Lemma B.6.
Let $G(x,y,\theta):\mathbb{R}^{2}\times\Theta\to\mathbb{R}$ be a function such that $G(x,\cdot,\theta)\in C^{2}(\mathbb{R})$ for each $x\in\mathbb{R}$ and $\theta\in\Theta$. Suppose that, for each $x\in\mathbb{R}$ and $k=1,2,\dots,l\ (l\geq 2)$, $\partial_{y}^{k}G(x,y,\theta)$ is of polynomial growth w.r.t. $y$ uniformly in $\theta\in\Theta$, $G(x,x,\theta)=0$, and that there exists a constant $M>0$ such that $|\partial_{y}^{l+1}G(x,y,\theta)|\leq M$ for all $x,y\in\mathbb{R}$ and $\theta\in\Theta$. Then it holds that
$\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}G(Y^{n}_{i-1},Y_{i}^{n},\theta)$ $\stackrel{{\scriptstyle p}}{{\longrightarrow}}\frac{\partial^{2}K(0)}{2K(0)}\int_{\mathbb{R}}\left[\partial_{y}G(z,z)\,z-\partial_{y}^{2}G(z,z)K(0)\right]\,\phi_{K(0)}(z)\,\mathrm{d}z,$
uniformly in $\theta\in\overline{\Theta}$ under $h_{n}\to 0$ as $n\to\infty$, where $Y_{i}^{n}=X_{t_{i}}-\int_{0}^{t_{i}}\mu_{\widehat{\xi}_{n}}(s)\,\mathrm{d}s$ and $\widehat{\xi}_{n}$ is a consistent estimator for $\xi_{0}$: $\widehat{\xi}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\xi_{0}$.
Proof.
Let us fix $\theta=(\xi,\sigma)\in\overline{\Theta}$. Define
$T_{n}(\theta):=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}G\left(Y_{i-1}^{n},Y_{i}^{n},\theta\right),$
where $Y_{i}^{n}:=X_{t_{i}}-\int_{0}^{t_{i}}\mu_{\widehat{\xi}_{n}}(s)\,\mathrm{d}s$ with a consistent estimator $\widehat{\xi}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\xi_{0}$.
By the model definition $X_{t}=Z_{t}+\int_{0}^{t}\mu_{\xi_{0}}(s)\,\mathrm{d}s$, we have:
$X_{t_{i}}-X_{t_{i-1}}=Z_{t_{i}}-Z_{t_{i-1}}+\int_{t_{i-1}}^{t_{i}}\mu_{\xi_{0}}(s)\,\mathrm{d}s.$
Thus,
$Y_{i}^{n}-Y_{i-1}^{n}=(Z_{t_{i}}-Z_{t_{i-1}})+r_{i}^{n},\quad\text{where }r_{i}^{n}:=\int_{t_{i-1}}^{t_{i}}(\mu_{\xi_{0}}(s)-\mu_{\widehat{\xi}_{n}}(s))\,\mathrm{d}s.$
Using Taylor expansion of $y\mapsto G(x,y,\theta)$ around $y=x$, we get
$G(x,y,\theta)$ $=\partial_{y}G(x,x,\theta)(y-x)+\frac{1}{2}\partial_{y}^{2}G(x,x,\theta)(y-x)^{2}+R(x,y,\theta),$
with the remainder $R(x,y,\theta)=\frac{1}{6}\partial_{y}^{3}G(x,\zeta,\theta)(y-x)^{3}$ for some $\zeta$ between $x$ and $y$. Then,
$T_{n}(\theta)$ $=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\Bigg{\{}\partial_{y}G(Y_{i-1}^{n},Y_{i-1}^{n},\theta)(Y_{i}^{n}-Y_{i-1}^{n})+\frac{1}{2}\partial_{y}^{2}G(Y_{i-1}^{n},Y_{i-1}^{n},\theta)(Y_{i}^{n}-Y_{i-1}^{n})^{2}+R_{i}^{n}(\theta)\Bigg{\}}$
$=:A_{n}(\theta)+B_{n}(\theta)+C_{n}(\theta).$
with $R_{i}^{n}(\theta)$ as above. We now evaluate $A_{n}(\theta)$ in detail:
$A_{n}(\theta)$ $=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\partial_{y}G(Y_{i-1}^{n},Y_{i-1}^{n},\theta)(\Delta_{i}^{n}Z+r_{i}^{n})$
$=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}},\theta)\Delta_{i}^{n}Z+R^{\prime}_{n}(\theta),$
where the error term $R^{\prime}_{n}(\theta)$ is decomposed as
$R^{\prime}_{n}(\theta)$ $=\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\left[\partial_{y}G(Y_{i-1}^{n},Y_{i-1}^{n},\theta)-\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}},\theta)\right]\Delta_{i}^{n}Z$
$\quad+\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\partial_{y}G(Y_{i-1}^{n},Y_{i-1}^{n},\theta)r_{i}^{n}.$
For the first term, using the integral form of the mean value theorem, we write:
$\partial_{y}G(Y_{i-1}^{n},Y_{i-1}^{n},\theta)-\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}},\theta)=\int_{0}^{1}\nabla_{1}\partial_{y}G(Z_{t_{i-1}}+u\delta_{i-1}^{n},Z_{t_{i-1}}+u\delta_{i-1}^{n},\theta)\cdot\delta_{i-1}^{n}\,\mathrm{d}u,$
where $\delta_{i}^{n}=\int_{0}^{t_{i}}[\mu_{\xi_{0}}(s)-\mu_{\widehat{\xi}_{n}}(s)]\,\mathrm{d}s$, so that the difference is bounded as
$\left|\partial_{y}G(Y_{i-1}^{n},Y_{i-1}^{n},\theta)-\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}},\theta)\right|\leq C(1+|Z_{t_{i-1}}|^{k})|\delta_{i-1}^{n}|,$
for some constant $C$ and $k\geq 0$, using the polynomial growth of $\nabla_{1}\partial_{y}G$. By Lemma A.5, this yields
$\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\left|\partial_{y}G(Y_{i-1}^{n},Y_{i-1}^{n},\theta)-\partial_{y}G(Z_{t_{i-1}},Z_{t_{i-1}},\theta)\right||\Delta_{i}^{n}Z|=O_{p}(h_{n})\to 0.$
The second term is bounded by
$\left|\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\partial_{y}G(Y_{i-1}^{n},Y_{i-1}^{n},\theta)r_{i}^{n}\right|\leq\frac{C}{nh_{n}}\sum_{i=1}^{n}(1+|Z_{t_{i-1}}|^{k})o_{p}(h_{n})=o_{p}(1),$
again using Lemma A.5 and the consistency $\widehat{\xi}_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\xi_{0}$. Therefore, using Lemma A.5 again, we conclude that
$A_{n}(\theta)\stackrel{{\scriptstyle p}}{{\longrightarrow}}-\frac{\partial^{2}K(0)}{2K(0)}\int_{\mathbb{R}}\partial_{y}G(z,z,\theta)z\phi_{K(0)}(z)\,\mathrm{d}z,$
The term $B_{n}(\theta)$ can be evaluated similarly using ergodic theory and standard limit theorems.
As for $C_{n}(\theta)$, recall that
$R_{i}^{n}(\theta)=\frac{1}{6}\partial_{y}^{3}G\left(X_{t_{i-1}},\zeta_{i},\theta\right)(X_{t_{i}}-X_{t_{i-1}})^{3}.$
By the assumption that $|\partial_{y}^{3}G(x,y,\theta)|\leq M$ and the fact that $X_{t_{i}}-X_{t_{i-1}}=O_{p}(h_{n}^{1/2})$, we get
$|C_{n}(\theta)|$ $\leq\frac{M}{6nh_{n}^{2}}\sum_{i=1}^{n}|X_{t_{i}}-X_{t_{i-1}}|^{3}$
$=O_{p}\left(\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}h_{n}^{3/2}\right)=O_{p}(h_{n}^{1/2})\to 0.$
Hence, the result follows. ∎
Remark B.2.
When G(x,y)=f(yx)G(x,y)=f(y-x) for a function $f\in C^{2}(\mathbb{R})$ with polynomial growth, the limit in Lemma B.6 simplifies to
$\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}f(Y^{n}_{i}-Y^{n}_{i-1})\to^{\mathbb{P}}-\frac{1}{2}\,\partial^{2}K(0)\cdot f^{\prime\prime}(0),$
as $n\to\infty$, where $Y^{n}_{i}=X_{t_{i}}-\int_{0}^{t_{i}}\mu_{\widehat{\xi}_{n}}(s)\,\mathrm{d}s$ and $\widehat{\xi}_{n}\to^{\mathbb{P}}\xi_{0}$. This result is consistent with Lemma B.1, which considers functions of the normalized increment
$\frac{\Delta_{i}^{n}X-\mu_{\xi}(t_{i-1})h_{n}}{h_{n}}\approx\frac{Y^{n}_{i}-Y^{n}_{i-1}}{h_{n}},$
and shows convergence of the empirical average of $f$ evaluated at that normalized quantity to the expectation under the normal distribution with mean zero and variance $\partial_{t}^{2}K(0)$. In this case, the asymptotic mean reduces to a constant multiple of $f^{\prime\prime}(0)$, matching the second moment of the limiting distribution.
Remark B.3.
Unlike Lemma B.1, the convergence in Lemma B.6 does not generally hold uniformly with respect to $\xi\in\Xi$. This difference arises from the nature of the function $G(x,y,\theta)$. In Lemma B.1, the function depends on only the factor (xy)(x-y), and the integral term $\int_{t_{i-1}}^{t_{i}}\mu_{\xi}(s)\,\mathrm{d}s$ is of order $O(h_{n})$ uniformly in $\xi$, allowing the substitution of $\xi$ without affecting the asymptotics. In contrast, the general form of $G(Y^{n}_{i-1},Y^{n}_{i},\theta)$ in Lemma B.6 does not necessarily exhibit such cancellation or stability, and thus the impact of the drift term cannot be neglected. Therefore, it is necessary to incorporate the consistent estimator $\widehat{\xi}_{n}$ to appropriately correct for the drift component.
Appendix C Ergodicity for Gaussian processes
C.1 Fundamental ergodic theorems
  • Let $Z$ be a (continuous) Gaussian process on a canonical space $(\Omega,\mathcal{F},{\mathbb{P}})$, where $\Omega=C(\mathbb{R})$. (So $\mathbb{P}$} is the distribution of $X$)
  • $\theta_{\tau}:\Omega\to\Omega$: the shift operator such that, for each $\omega\in\Omega$,
    $\theta_{\tau}\omega(t)=\omega(t-\tau),\quad\tau,t\in\mathbb{R}.$
Theorem C.1 (Maruyama [14] or Krishnapur [13]).
Let $Z$ be a centered stationary Gaussian process on $\mathbb{R}^{d}$ with continuous covariance kernel $K$ and the spectral measure $\mu$: for any $t,s\in\mathbb{R}$,
$\mathbb{E}[Z_{t}]=0;\quad\mathbb{E}[Z_{t}Z_{s}]=K(t-s);\quad K(h)=\int_{\mathbb{R}}e^{ihx}\mu(\mathrm{d}x).$
Then
  • (i)
    $Z$ is ergodic if and only if $\mu$ has no atom.
  • (ii)
    $Z$ is weakly mixing if and only if $K(t)=o(1)$ as $|t|\to\infty$.
Theorem C.2 (Birkoff’s ergodic theorem).
$Z$ is ergodic if and only if, for any $g\in L^{1}({\mathbb{P}})$,
$\frac{1}{T}\int_{0}^{T}g(\theta_{\tau}\omega)\,\mathrm{d}\tau\to\mathbb{E}[g],\quad T\to\infty,$
almost surely or $L^{1}({\mathbb{P}})$ sense. Therefore, it converges at least in probability.
C.2 Utility format in applications
  • Suppose that $Z=(Z_{t})_{t\geq 0}\sim GP(0,K)$ with $K(t)\to 0$ as $t\to\infty$.
  • Let $\pi_{t}\ (t\geq 0)$ is a canonical projection: $\pi_{t}Z=Z_{t}$.
For a function $f:\mathbb{R}\to\mathbb{R}$, of polynomial growth: $|f(x)|\lesssim(1+|x|)^{C}$, and a fixed $t\geq 0$, put $g(\omega)=f\circ\pi_{t}(\omega)\ (\omega\in\Omega)$, then
$g\sim^{d}f(Z_{t})=:f(G),\quad G\sim{\cal N}(0,K(0)),$
by the stationarity, and Thereom C.2 says that
$\frac{1}{T}\int_{0}^{T}g(X(\cdot-\tau))\,\mathrm{d}\tau$ $=\frac{1}{T}\int_{0}^{T}f(Z_{T-\tau})\,\mathrm{d}\tau=\frac{1}{T}\int_{0}^{T}f(Z_{u})\,\mathrm{d}u$
$\stackrel{{\scriptstyle p}}{{\longrightarrow}}\mathbb{E}[f(G)]=\int_{\mathbb{R}}f(z)\phi_{K(0)}(z)\,\mathrm{d}z,\quad T\to\infty,$
where $\phi_{\mu,\Sigma}(z)$ is the probability density of ${\cal N}(\mu,\Sigma)$.
Corollary C.1.
Let $Z=(Z_{t})_{t\geq 0}$ be a centered stationary Gaussian process with $K(t)\to 0\ (t\to\infty)$. Then it holds for any function $f:\mathbb{R}\to\mathbb{R}$, of polynomial growth that
$\frac{1}{T}\int_{0}^{T}f(Z_{u})\,\mathrm{d}u\to\mathbb{E}[f(Z_{0})]\quad\mbox{$as$ or in $L^{1}$},\quad T\to\infty,$ (C.1)
for any measurable function $f:\mathbb{R}\to\mathbb{R}$ such that $\mathbb{E}[f(Z_{0})]<\infty$.